### v1

# Irreversibility and symmetry breaking in the creation and annihilation of defects in active living matter

Abstract
Active living matter continuously creates and annihilates topological defects in a process that remains poorly understood. Here, we investigate these dynamics in two distinct active living systems—swarming bacteria and human bronchial epithelial cells. Despite their entirely different evolutionary origins, biological functions, and physical scales, both systems exhibit half-integer defects, consistent with the nematic phase. However, in contrast to active nematic theory, we find that defect creation and annihilation undergoes spatial symmetry breaking. We propose that this results from a fundamental dualism between nematic structural organization and generated polar forces, which are intrinsic to living systems. Furthermore, our estimation of entropy production reveals that creation and annihilation are not reversed processes. Our findings challenge conventional nematic models and emphasize the role of defect-mediated dynamics in non-equilibrium biological systems as a major source of entropy production.
I Introduction
Active nematic materials are typically comprised of elongated components that generate stress at the microscopic scale, leading to large-scale motion Sanchez et al. (2012); Guillamat et al. (2018); Tan et al. (2019); Kumar et al. (2018). Unlike passive systems, active matter remains in a state of perpetual motion and reorganization, which is driven by internally generated forces Sokolov et al. (2010); Beta et al. (2023). A hallmark of active nematic materials is the continuous creation and annihilation of topological defects Marchetti et al. (2013), which result from a balance between active and elastic forces Giomi (2015); Zhou et al. (2014); Serra et al. (2023); Guillamat Bassedas et al. (2017). Of particular interest are living systems, such as bacterial collectives Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017) and tissue cells Saw et al. (2017); Balasubramaniam et al. (2021); Blanch-Mercader et al. (2018), because their physical properties are also related to their biological functions Be’er and Ariel (2019); Aranson (2022); Doostmohammadi and Ladoux (2022). In living matter, active stress originates from metabolism and the consumption of chemical fuel, which enables bacteria and cells to move and generate forces on their neighbors. Elasticity arises from an elongated cellular shape, leading to liquid-crystal-like properties Be’er and Ariel (2019); Aranson (2022); Doostmohammadi and Ladoux (2022). The resulting defects are mesoscopic objects involving tens to hundreds of cells and governed by intrinsic spatiotemporal scales that are independent of system size Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017); Saw et al. (2017); Balasubramaniam et al. (2021); Blanch-Mercader et al. (2018).
Experiments with bacterial colonies and cell monolayers consistently reveal half-integer defects, in accordance with nematic symmetry Li et al. (2019); Yashunsky et al. (2024); Shimaya and Takeuchi (2022); Copenhagen et al. (2021); Genkin et al. (2017); Duclos et al. (2017); Blanch-Mercader et al. (2018); Saw et al. (2017); Kawaguchi et al. (2017); Balasubramaniam et al. (2021); Ienaga et al. (2023). Recent studies suggested that nematic symmetry can also manifest in systems with directionality Venkatesh et al. (2025), such as models of self-propelled rods Shi and Chaté (2018); Großmann et al. (2020) and bacterial experiments Head et al. (2024). Intriguingly, some nematically ordered systems display polar characteristics under specific conditions, indicating the coexistence of nematic and polar symmetries Huber et al. (2018); Meacock et al. (2021); Maroudas-Sacks et al. (2021); Amiri et al. (2022); Lacroix et al. (2024). This interplay of symmetries is fundamental to understanding biological processes, such as biofilm formation and the organization of cellular tissues, yet remains underexplored experimentally.
Defect interactions have been primarily investigated through numerical simulations Thampi et al. (2014); Shankar et al. (2018); Patelli et al. (2019); Bonn et al. (2022); Pearce et al. (2021). Elastic forces between defects predict Coulomb-like interactions, as well as additional forces and torques that rely on their relative orientation Shankar et al. (2022); Pearce and Kruse (2021); Tang and Selinger (2017); Vromans and Giomi (2016). It has been suggested that nematic active stresses also lead to the generation of hydrodynamic flows around the defects which results in self-propulsion of $^{1}/_{2}$ defects. Rotational diffusion can reduce persistence in $^{1}/_{2}$ defect motion, preventing unbinding from $^{1}/_{2}$ counterparts, and affecting defect density in steady states Shankar et al. (2018). However, much of the theoretical framework remains experimentally unverified.
Recent experiments on microtubule-based active nematics and swarming bacteria have uncovered a broad spectrum of defect organization, ranging from giant number fluctuations to hyperuniformity Yashunsky et al. (2024); Fernandez-Nieves et al. (2024); Pearce et al. (2021). In fibrosarcoma cell cultures, $^{1}/_{2}$ defects align toward the edge of the colony, generating chiral edge flows through local chiral active stress Yashunsky et al. (2022). Spontaneous mirror (chiral) symmetry breaking has been demonstrated for $^{1}/_{2}$ defects, which specifically concentrate at the boundary between vorticity-dominated and strain-rate-dominated regions in microtubule-kinesin films, highlighting their strong coupling to large-scale hydrodynamic fluxes within the material Head et al. (2024). However, the creation and annihilation processes, which we identified below as primary sources of energy injection and dissipation in living systems, are underexplored and necessitating experimental validation of current theoretical predictions.
In this paper, we experimentally study swarming bacteria Be’er and Ariel (2019) and human bronchial epithelial cells (HBECs) Blanch-Mercader et al. (2018)—distinct systems that both display nematic half-integer defects. Our findings reveal that the creation and annihilation of defects follow distinct, previously unrecognized trajectories that spontaneously break mirror symmetry—a feature absent in standard active nematic models. We also find that these trajectories exhibit time-reversal symmetry breaking, a hallmark of non-equilibrium processes. Statistical analysis reveals that defect creation and annihilation drive substantial entropy production, underscoring their irreversible nature and critical role in active matter organization. Finally, we demonstrate why these observations can be explained by taking into account directional (polar) self-propulsion in bacteria and cells.
Figure 1: Director field and nematic defects in swarming Bacillus subtilis and human bronchial epithelial cells (HBECs) monolayers. The nematic director field, colored according to the orientations (ranging from $\pi/2$ to $\pi/2$), and half-integer defects in (a) swarming B. subtilis and (b) HBECs. $^{1}/_{2}$ types are represented as pink comet shapes and $^{1}/_{2}$ as violet tripods. (c, d) Rates of pair-creation (blue) and annihilation (red) of oppositely charged $\pm^{1}/_{2}$ defects; (c) swarming B. subtilis and (d) HBECs.
II Results
Half-integer defects in bacteria and cells
We study the creation and annihilation of defects in two highly active, quasi-turbulent living systems of swarming Bacillus subtilis bacteria Yashunsky et al. (2024) and human bronchial epithelial cells (HBECs) Blanch-Mercader et al. (2018). Despite the extreme disparity in the time scales and dimension (Fig. 1), both systems share fundamental characteristics of active extensile nematics Giomi (2015); Mart´ınez-Prat et al. (2021): (i) A steady-state mixture of half-integer defects (Figs. 1a-b), which is obtained through continuing creation and annihilation events (Figs. 1c-d), (ii) a counter-rotating vortex pattern around defects with (iii) an exponential distribution of vortex sizes Yashunsky et al. (2022); Blanch-Mercader et al. (2018).
The defects are created and annihilated in pairs with opposite charges. In both systems, the rates of creation and annihilation events fluctuate over time but stabilizes to a steady state (Figs. 1c-d), implying an approximately constant number of defects over time with a total charge that fluctuates around zero. We characterize the motion of defect pairs through these observables, see the sketch in Fig. 2.
  • $t$: time from creation or annihilation event (negative time indicates time before annihilation).
  • $d$: the distance between the $^{1}/_{2}$ and $^{1}/_{2}$ defects in the pair.
  • $\gamma$: the argument of the vector from the $^{1}/_{2}$ to the $^{1}/_{2}$ defect core, rotated so that $\gamma(t=0)=0$.
  • $\varphi^{+}$: the angle between the orientation of the “tail” of the comet-shaped $^{1}/_{2}$ defect and the vector connecting the pair. Thus, $\varphi^{+}=0$ indicates that the tail is pointing away from the $^{1}/_{2}$ defect.
  • $\varphi^{-}$: the angle between the orientation of one of the $^{1}/_{2}$ legs and the line connecting the pair. We use the smallest angle out of the three legs. For example, $\varphi^{-}=0$ implies that one of the legs is pointing towards the $^{1}/_{2}$ defect.
  • $\delta=\varphi^{+}-3\varphi^{-}-\pi$: pair orientation offset. This gives a measure of how compatible the orientations of the defects are with each other. $\delta=0$ suggests the defects are at the orientations that minimize the variation in the nematic field between themPearce and Kruse (2021).
Figure 2: Configuration of $^{1}/_{2}$ and $^{1}/_{2}$ defect-pairs during creation and annihilation. Sketch: Definitions of orientational angles $\varphi^{+}$ (pink) and $\varphi^{-}$ (violet) for paired $^{1}/_{2}$ and $^{1}/_{2}$ defects relative to the pair axis connecting between $^{1}/_{2}$ and $^{1}/_{2}$ defect centers (horizontal black line). (a, b) Distribution of the pair orientation offset angle, $\delta=\varphi^{+}-3\varphi^{-}-\pi$, as a function of time for creation (blue) and annihilation (red) events. $\delta=0$ implies that the nematic director fields of the paired-defects are synchronized. (c, d) Distribution of $\varphi^{+}$ as a function of time for the creation (blue) and annihilation (red) events. The creation/annihilation instant is indicated by $t=0$, with negative time indicating time to annihilation, and positive time is from creation. Red and blue color intensities indicate probability density. (e-g) Alignment of defect pairs, classified by the mean orientation of the $^{1}/_{2}$ defect during its trajectory, $\langle\varphi^{+}\rangle$. Angular distributions of $\varphi^{+}(d<r_{0})$ and $\varphi^{-}(d<r_{0})$ within $r_{0}=15~\mu m$ for bacteria and $35~\mu m$ for cells are shown for “up” ($\langle\varphi^{+}\rangle>0$, green) and “down” ($\langle\varphi^{+}\rangle<0$, yellow) configurations in bacteria (e, f) and cells (g, h).
Spontaneous breaking of the mirror symmetry in defect orientation
At annihilation or creation events, defect pairs are in phase ($\delta=0$; see Figs. 2a-b), indicating they are perfectly aligned. When $\delta~\neq 0$, the elastic energy between the defects diverges at small separations, indicating that the defects align at very short distances Pearce and Kruse (2021).
We next look at the alignment of the $^{1}/_{2}$ defect with the line connecting the defects in the pair, captured by the parameter $\varphi^{+}$. We see a bimodal distribution for $\varphi^{+}$ immediately before annihilation with peaks at $\pm\pi/2$; see Figs. 2c-d. This indicates that the tail of the positive defect is perpendicular to the direction of annihilation at the end of this process. A similar but less pronounced effect is also visible immediately after creation. This bimodality indicates a spontaneously broken chiral symmetry in the orientation of the defects. To quantify this effect, we split all trajectories into two groups based on their time-averaged orientation across the full trajectory, either “up” (average $\varphi^{+}>0$) or “down” (average $\varphi^{+}<0$). The orientation $\varphi^{-}$ follows a similar symmetry breaking, consistent with $\delta=0$ (Figs. 2e-f).
Figure 3: Spiraling trajectories of creation and annihilation. Sketch: Co-rotation angle $\gamma$—the line connecting between the location of defect centers with respect to the angle close to creation/annihilation instant ($\gamma_{d\rightarrow 0}$). (a, b) Distance between $\pm^{1}/_{2}$ defect-pairs as a function of time after creation (blue) and before annihilation (red), for bacteria and cells. The creation/annihilation instant is indicated by $t({d\rightarrow 0})$, with negative values indicating time to annihilation, and positive ones indicating time from creation. Solid lines represent the average trajectories, while red and blue color intensities indicate probability density. (c, d) Time evolution of $\gamma$ (the angle of the line connecting the defect-pairs, aligned to $\gamma_{d\rightarrow 0}=0$). Solid lines represent averages for “up” (green) and “down” (yellow) configurations. Shaded areas indicate the standard deviation. (e, f) Average trajectories of $^{1}/_{2}$ (pink) and $^{1}/_{2}$ (violet) defects during creation and annihilation, measured in the lab frame of reference. Trajectories are centered around the creation/annihilation position (black point), rotated to $\gamma_{d\rightarrow 0}$ and mirrored to the “up” configuration.
Spiraling trajectories of creation and annihilation
The separation rate between paired defects (Figs. 3a-b) is characterized by increased speed close to the creation or annihilation events, consistent with theoretical predictions for defect annihilation in passive systems Pearce and Kruse (2021). However, defects do not move in a straight line. This is quantified by the orientation of the vector connecting the defects changes over time, $\gamma(t)$. Generally, $\gamma$ varies monotonically with time, indicating that defect pairs tend to co-rotate during the process of annihilation or creation. Furthermore, the chirality of the co-rotation is correlated with the “up” or “down” configuration (Figs. 3c-d).
Lastly, Figs. 3e-f show averaged trajectories for creation and annihilation events in the “up” configuration. This reflects the spiraling motion toward (away) each other during annihilation (creation). These trajectories are similar to those predicted for the annihilation of topological defects with an initial phase disparity Pearce and Kruse (2021).
Figure 4: Ensemble-averaged flow fields near $\pm^{1}/_{2}$ defect pairs. (a, b) Average velocity fields (black arrows) and vorticity distributions (heatmap) around defects. (a-d) Flow fields around paired $^{1}/_{2}$ and $^{1}/_{2}$ defects, averaged separately for the “up” (a, b) and “down” (a, d) configurations. e, f) Flow fields around unpaired $^{1}/_{2}$ and $^{1}/_{2}$ defects. For clarity, only results for bacteria close to annihilation are shown. The angular orientation of defects is illustrated in the schematic. Arrow and spatial scales are indicated. Scale bar: $10~\mu m$. The number of fields averaged to obtain each panel is: (a) $5459$; (b) $5391$; (c) $5336$; (d) $5323$; (e) $89027$; (f) $88267$.
Spontaneous breaking of the mirror symmetry in flow fields
We find that flows around defect pairs also exhibit mirror symmetry breaking (Figs. 4a-d, Extended Data Figs. 1-2). Flows were measured separately for the “up” and “down” orientations. In bacteria, the forward flow at the core of the defect $^{1}/_{2}$ is offset relative to the defect symmetry axis. The direction of the tilt changes between annihilation and creation (Extended Data Fig. 1). In particular, for both bacteria and cells, the flow at the core of the $^{1}/_{2}$ defects exhibits a distinct rotational behavior, rotating clockwise for “up”-oriented configurations (Fig. 4B, Extended Data Figs. 1b,f and 2b,f) and counterclockwise for “down”-oriented configurations (Fig. 4d, Extended Data Figs. 1d,h and 2d,h). The flows around unpaired defects (Figs. 4e-f) retain mirror symmetry, in agreement with theoretically predicted flows around isolated defects in extensile active nematics.
Figure 5: The interaction between nematic and polar symmetries. (a) Defect creation via bend instability with the overlaid polar field, indicated by red arrows; a grain boundary in the polar field is introduced indicated by the blue dashed line. The polar field leads to a torque on the defects, indicated by the green arrow. The torque will reorient the $^{1}/_{2}$ defect, causing it to self-propel around the $^{1}/_{2}$ defect. Shown here is the “up” configuration. (b) Nematic and polar fields around a pair of defects before annihilation. Here the grain boundaries do not directly connect. The same torque is present around the $^{1}/_{2}$ defect that can lead to a spiral trajectory. (c) & (d) Predicted Stokes flow field around a (c) $^{1}/_{2}$ and (d) $^{1}/_{2}$ defect with additional active force along the polar direction; shown in the “up” configuration.
The source of mirror symmetry breaking
In addition to nematic order, bacteria and cells also exhibit a polar symmetry due to their directed self-propulsion mechanism. Combining nematic interactions with active stresses of various symmetries can significantly alter active nematic behavior Ardaševa et al. (2025); Großmann et al. (2020).
To describe cell self-propulsion, we introduce a polar field, $\underline{p}$, indicating the direction of cell movement relative to the substrate, alongside the nematic field that describes the cells’ long axes. This can be visualized as two superimposed parallel fields: one with lines for the long axes and another with arrows for the head direction (see Fig. 5a). For simplicity, we assume that both $\underline{p}$ and $\underline{p}$, the latter pointing in the opposite direction, are compatible with the same nematic field.
During creation in extensile nematics, defects arise from a bend instability that introduces a net curl to the polar field and breaks chiral symmetry (Fig. 5a), with a particularly strong net curl around the $^{1}/_{2}$ defect (green arrow; Fig. 5a). Notably, the mirror image of this process produces defects with opposite chirality, indicating that defect chirality selection is driven by spontaneous symmetry breaking. In this context, distinct “up” and “down” defect configurations should emerge during creation.
Around an isolated defect, the polar field exhibits a net polarization and curl that depend on the grain boundary position, see SI for a derivation. This polar field indicates the force direction exerted by neighboring particles, generating both force and torque on the defect core (Fig.  5a). A similar symmetry breaking mechanism has been noted in simulations of self-propelling thin rods and flexible polymers Palmer et al. (2022); Prathyusha et al. (2018); Joshi et al. (2019).
We suggest that defect dynamics and flows around the defects in our experiments result from a combination of torque and extensile active stress. For a $^{1}/_{2}$ defect, this leads to an offset force along the defect’s symmetry axis and an additional net torque. $^{1}/_{2}$ defect experiences smaller net torques and forces, allowing for self-propulsion and rotation. The calculated Stokes flow field around isolated defects (Figs. 5c-d) indicates that immediately after the defects are created, flow patterns exhibiting opposite chiral symmetry are equally likely.
This symmetry breaking can explain the distinct behavior of “up” and “down” configurations as follows. Defects are generated through a bend instability induced by extensile active stress, hence are generated with $\varphi^{+}\approx 0$. After creation, $^{1}/_{2}$ move in a curved trajectory, spiraling away from its $^{1}/_{2}$ pair due to the non-zero curl resulting from the polar field. Figure 5a illustrates a situation where a $^{1}/_{2}$ defect moves, rotating clockwise. Consequently, this leads to the defect pair to also rotate clockwise, which we refer to as “up” configuration (see Fig. 4).
The polar grain boundary connected to the defect core is topologically constrained and persists throughout the defect’s lifetime. As a result, the same broken symmetry influences the defect annihilation process, as illustrated in Fig. 5b. When a $^{1}/_{2}$ defect approaches annihilation, two key forces dictate its trajectory. The first is the elastic attraction to the $^{1}/_{2}$ defect with which it is annihilating, and the second is the active self-propulsion force. The chiral torque acting on the defect causes it to rotate, so it does not self-propel directly toward the $^{1}/_{2}$ defect; however, the elastic forces prevent complete escape. Consequently, the resulting trajectory forms an inward spiral toward annihilation, as depicted in Figs. 3e-f. In contrast to the creation process, annihilation is independent of $\varphi^{+}$ (assuming $\delta=0$). Thus, defects can annihilate even when $\varphi^{+}\neq 0$, as shown in Figs. 2c-d. Additionally, the reverse of this process is equally likely, resulting in the observed “up” and “down” configurations during annihilation.
Time-reversal symmetry breaking
Active systems are typically irreversible, which is manifested in strictly positive entropy production (EP). Loosly speaking, EP is defined as the Kullback-Leibler divergence (KLD) from the distribution of trajectories compared to the distribution of time-reversed trajectories Seifert (2012); Fodor et al. (2022). In some systems, such as Langevin particles satisfying the Einstein relation, EP equals (in units of temperature) to the dissipated heat Fodor et al. (2022); Ro et al. (2022). It has been suggested that EP also plays a role in active systems, and is indicative of the amount of extractable work Fodor et al. (2022); Ro et al. (2022). Previous attempts to quantify EP in bacterial and cellular dynamics proved challenging. In Ro et al. (2022), local EP in swimming bacteria was estimated using a compression-based methods, and found positive EP mainly close to walls or barriers. However, negligible or no EP was observed in the bulk, despite the fact that the system is highly active and out of equilibrium.
Obtaining accurate statistics of the joint distribution of cellular or bacterial trajectories is challenging because of the high dimensionality of such a dataset as estimating KLD suffers from severe undersampling of the phase space Ro et al. (2022). Here, defects provide physically motivated coarse-graining of the micro-scale motion into a mesoscopically observable feature. As a consequence, all the estimates reported here are lower bounds to the actual EP.
In extensile-active nematics, $^{1}/_{2}$ defects move in a direction that is opposite to the “comet-tail” Giomi (2015). To quantify this effect, Table 3 shows the average projection of the defect direction onto the velocity, which is indeed negative. We can calculate a similar measure by contracting the velocity of a $^{1}/_{2}$ defect with the third rank tensor that describes its orientation (see Methods). Interestingly, we find a non-zero average value (Table 3) that is not expected in a typical active nematic. This is consistent with the net force we calculate on the core of a $^{1}/_{2}$ defect resulting from the polar active forcing (see SI).
Table 3 also shows results for EP estimates. For bacteria, we find a high value of EP in $^{1}/_{2}$ defects, and lower (but statistically significant) for $^{1}/_{2}$ ones. For cells, the EP in $^{1}/_{2}$ cells is an order of magnitude smaller than for bacteria. For $^{1}/_{2}$ it is below the error range.
Next, we turn to analyze creation and annihilation events. Figures 2c-d show that the distribution of $^{1}/_{2}$ is different close to creation compared to annihilation. Similar broken symmetries can also be observed in the average trajectories (Figs. 3e-f) and the flow fields (Extended Data Figs. 1 and 2). To quantify this, Table 2 details the results for EP estimates. The dimension $D$ indicated the length of the trajectories used (starting at the event). Indeed, our results show positive EP both for bacteria and cells.
Table 1: Entropy production (EP) estimation for isolated defects. The row noted projection shows the projection of the defect orientation on the velocity. The row noted EP shows a statistical estimate for a lower bound for the entropy production rate obtained for defect trajectories. Values show the estimate $\pm$ the standard error. Since this is a lower bound, a value that is statistically larger than zero implies non reversible dynamics. See the methods section for details.
Bacteria Cells
Defect charge $^{1}/_{2}$ $^{1}/_{2}$ $^{1}/_{2}$ $^{1}/_{2}$
projection $2\pm 0.003 0.0$$05\pm 0.003 0.0$$8\pm 0.003 0.0$$01\pm 0.003$
EP $94\pm 0.03 0.0$$005\pm 0.003$ $18\pm 0.01 0.0$$0006\pm 0.0007$
Table 2: Entropy production (EP) estimation for defect pairs. The table shows different statistical estimates for a lower bound for the entropy production rate obtained for defect trajectories. The dimension $D$ indicated the length of the trajectories used (starting at the event). Values show the estimate $\pm$ the standard error. Since this is a lower bound, a value that is statistically larger than zero implies non reversible dynamics. See the methods section for details.
*Note that the KLD is a non-negative quantity. However, the estimator are random variable with an error that is normally distributed with the noted standard deviation. Hence, zero EP my result in a negative estimate that is several times the standard error. In statistical terms, the hypothesis that $EP=0$ cannot be negated.
Bacteria Cells
Event creation annihilation cross creation annihilation cross
$EP(D=2)$ $075\pm 0.008 0.0$$4\pm 0.008 0.0$$01\pm 0.005^{*}$ $0.63\pm 0.03$ $0.55\pm 0.03$ $0.4\pm 0.3$
$EP(D=5)$ $0.11\pm 0.01$ $14\pm 0.02 0.0$$06\pm 0.02$ $0.46\pm 0.04$ $47\pm 0.05 0.0$$04\pm 0.03^{*}$
Focusing on time-reversal, running the experimental movies backward, creation will appear as annihilation and vice versa. To this end, Table 2 shows the cross entropy between forward-creation/backward-annihilation and forward-annihilation/backward-creation. The dimension $D$ indicated the length of the trajectories used (starting at the event). The positive cross-entropy indicated that creation and annihilation are not time-symmetric, both for bacteria and cells.
Comparing the total EP rates, we need to consider the rate of creation and annihilation events. For bacteria, the time between events is about 1 sec (50 frames), hence the total EP due to movement is about 50 times larger than due creation-annihilation. For cells, the time between creation and annihilation is about 2 hrs (24 frames). Hence, EP due to movement is comparable to EP due to creation-annihilation. These differences are consistent with the observation that in bacteria, $^{1}/_{2}$ defects appear self-propelled in the sense that their trajectories are ballistic over short time periods Yashunsky et al. (2024), implying a relatively large EP. With cells, both $\pm^{1}/_{2}$ defects are diffusive Blanch-Mercader et al. (2018), their movement in the bulk is close to reversible, and EP is most apparent in creation and annihilation.
Conclusions
Living systems constantly produce and consume energy, keeping them far from thermodynamic equilibrium. This continuous energy input drives self-organization, often breaking spatial symmetries in ways that shape the biological function Bowick et al. (2022); Pisegna et al. (2024); Ardaševa et al. (2025). A compelling hypothesis suggests that living systems leverage topological mechanisms to regulate their organization, providing a unifying framework for understanding emergent behaviors across biological systems Shankar et al. (2022); Andersen et al. (2025).
Here, we explore this idea in two seemingly different systems: Bacillus subtilis bacterial swarms and human bronchial epithelial cells. Despite stark differences in biological function and physical properties, both systems exhibit active nematic organization, in which rod-shaped bacteria and elongated polygonal cells align and form dynamic topological defects. While bacteria propel in a thin fluid layer at low Reynolds numbers ($\sim 10^{-4}$) Be’er and Ariel (2019); Aranson (2022), HBECs move via substrate adhesion and traction forces. In both cases, motion is highly dissipative and sustained only through continuous energy consumption at the cellular level Vaidžiulytė et al. (2022). Both systems reveal nematic order, where $\pm^{1}/_{2}$ defects undergo continuing creation and annihilation events between defects pairs with opposite charge.
The non-equilibrium and irreversible nature of active living systems is unequivocal. Because bacteria and cells are self propelled, the dynamics is out of equilibrium even at the level of the individual. Therefore, the theoretical prediction is that irreversibility should be observed at the single-cell level. However, the difficulty in capturing irreversibility from individual positions and orientations is reasonable. For any observed single-cell trajectory, one can imagine a cell following the reversed path. Thus, EP from single cells is only expected to be captured by adding intra-cellular degrees of freedom corresponding to chemical and biological function, or possibly also the underlying self-propulsion mechanism.
We can think of defects as physically intuitive, virtual meso-scopic degrees of freedom. Virtual in the sense that they are not the particles of which the system is made of. Meso-scopic since each one is defined by a small local collective on several close cells. Our findings show that irreversibility is statistically evident locally at the level of the collective. Moreover, we find that defect creation and annihilation are not reciprocal processes that general entropy. This is also physically intuitive as continuous creation of defects in nematic systems is a hallmark of activity. In both experimental systems, entropy production during defect creation and annihilation accounts for a disproportionately large fraction of total entropy production. In cells, it is an order of magnitude higher, suggesting that these events are key drivers of non-equilibrium dynamics rather than incidental fluctuations. Still, one should remember that our estimates are merely lower bounds for the actual EP, which means that quantitative comparison is problematic.
We find that mirror symmetry, predicted by linear nematodynamic theory, spontaneously break in the process of creation and annihilation. Specifically, when close to each other, $^{1}/_{2}$ spontaneously align perpendicularly (“up” or “down”) to the pair axis (Fig. 2 e-g). Each of these configurations follows spiral-like trajectories rotating either in the clockwise (“up”) or counterclockwise (“down”) direction (Fig. 3 c-f), further indicating spontaneous breaking of the mirror symmetry.
We propose that directional self-propulsion disrupts the nematic symmetry that otherwise captures the coarse-grained behavior of bacterial swarms and epithelial cell monolayers.
This reinforces the role of topological defects in understanding principles of self-organization in living systems.
Our results challenge the standard active nematic description often applied to active living matter. By uncovering these universal physical principles, we take a step toward a deeper understanding of biological self-organization, revealing the fundamental role of symmetry in shaping life at multiple scales.
III Methods
Swarming bacteria
Experiments were conducted on Bacillus subtilis 3610, the “wild-type” (WT) strain, which is a rod-shaped flagellated species with dimensions $1~\mu m\times 7~\mu m$. Cells are stored at $80~^{\circ}C$ in frozen stocks. Fluorescently labeled variants, green and red, are used (amyE::PvegR0_sfGFP;specR and amyE::Pveg_R0_mKate;specR); labeling does not affect any known measured quantity. In order to resolve the cells in the crowded colony, we mix the two labeled strains at ratio $1:1$ (after separate overnight growth), which yields much sparser fields of view (see camera settings next). Swarm experiments are typically done in a standard Petri-dish (8.8cm8.8~cm in diameter), where the colonies are grown on soft agar ($0.5~\%$ and 25g/L25~g/L LB) and aged in ambient lab conditions ($24~^{\circ}C$ and $35~\%$ RH) for 4 days. The latter forms the thicker colony structure compared to results obtained in other past studies. Overnight cultures are grown from isolated colonies (which are grown from frozen stocks on hard agar plates ($2~\%$ and LB)) at shaking ($200$ rpm and $30~^{\circ}C$) for 18h18~h in LB liquid medium (2mL2~mL in a 15mL15~mL tube). A small drop of overnight culture ($4~\mu L$), either from a single strain, or a $1:1$ mix of the two labeled variants, is deposited on the agar in the middle of the plate. The colonies grow for a few hours in a $95~\%$ RH incubator at $30~^{\circ}C$. The swarm colonies form a quasi-2D structure with a thickness of $\sim 7~\mu m$, which is uniform along centimeter-wide distances. Observations are done using an optical microscope (Axio Imager Z2 Zeiss; 40×) operated at fluorescence mode; this magnification yields a $150~\mu m\times 150~\mu m$ observation frame. The system is equipped with a splitting device (Optosplit II) that enables dual excitation and acquisition. The two fluorescence fields are obtained for the same spatial field and time. The image is projected on a NEO camera with $1800\times 900$ pixels and at 50 frames per second so that the green cells are seen on the left panel while the red ones are seen on the right panel. In the current study we focus on the upper layer of the colony. The system operates with the dual fluorescence set Ex 59026x, beam splitter 69008bs, and Em 535/30; 632/60. A compensation lens (an integral part of the Optosplit II device) is used to adjust the focus in each of the panels so that both the green panel and the red one are in focus at the upper layer.
Human bronchial epithelial cells (HBECs)
Human bronchial epithelial cells (HBECs) cultured as previously reported. HBEC culture was executed in a supplemented keratinocyte serum-free medium with l-glutamine (Keratinocyte-SFM with l-glutamine; Gibco). Cells were maintained at $37~^{\circ}C$ under $5~\%$ $CO_{2}$ partial pressure and $95~\%$ relative humidity. For time-lapse experiments, cells were seeded at a density of approximately 640,000 cells per well in polystyrene (TPP) 12-well tissue culture plates. Time-lapse microscopy experiments were performed in phase contrast on an automated IX71 inverted microscope (Olympus) equipped with the same temperature, humidity, and $CO_{2}$ regulations as in culture (Life Imaging Services). Images were acquired using a $10\times$ objective, providing a $1.5~mm\times 1.5~mm$ observation frame. The intervals between images were set to 5 minutes.
Measurement of Velocity and Director Fields
Image analysis was conducted using custom Python code. To obtain the velocity field from microscopic images, we employed the Gunnar Farneback method available from OpenCV library producing displacement vectors for every pixel rather than a sparse set of points. The director field was computed using the second-moment matrix, implemented with the structure tensor function from the skimage Python library. Both the flow and director fields were computed with an averaging window size of $2.5~\mu m\times 2.5~\mu m$ for bacteria and $11.1~\mu m\times 11.1~\mu m$ for cells.
Defect Detection and Tracking
Nematic defects were identified using the winding number method, which quantifies the degree of rotation in the nematic field around any closed loop of nearest neighbors. When the loop encircles a defect, the winding number reflects the charge of that defect. The orientation of the defects was calculated following the method outlined by Vromans & Giomi (2016) Vromans and Giomi (2016). Angles are discrete, at integer multiples of $\pi/50$. Defect trajectories were automatically tracked using the TrackMate plugin in ImageJ software, which was combined with defect orientation data. This generated a list of defects along with their charges, angles, and locations.
Overall, the data on a single $^{1}/_{2}$ defect trajectory is a sequence of vectors $\textbf{p}_{0},...,\textbf{p}_{K}$, where each vector pi is in $R^{3}$, i.e., each frame consists of 3 scalar values, corresponding to the xyx-y coordinates and an angle, corresponding to the orientation of the $^{1}/_{2}$ defect. The data on a single $^{1}/_{2}$ defect trajectory is a sequence of vectors $\textbf{n}_{0},...,\textbf{n}_{K}$, where each vector $\textbf{n}_{i}$ is in $R^{5}$, i.e., each frame consists of 5 scalar values, corresponding to the x-y coordinates and the three angles, corresponding to the three directions of the $^{1}/_{2}$ defect. They may not be exactly $2\pi/3$ apart. Velocities were taken as the displacements between consecutive frames with no further smoothing.
Defect-Pair Detection
To identify defect-pairs, thresholds for distance ($\Delta d$) and time ($\Delta t$) were applied. The thresholds were set at $\Delta d=2~\mu m$ for bacteria and $\Delta d=11.1~\mu m$ for cells, with $\Delta t=0.04$ seconds for bacteria and 10 minutes for cells. Two oppositely charged defects, $a(t_{a},~x_{a},~y_{a})$ and $b(t_{b},~x_{b},~y_{b})$ were considered a created pair if their initial detection times, $t_{a}$ and $t_{b}$, satisfied $|t_{a}-t_{b}|<\Delta t$, and their initial locations, $x_{a},~y_{a}$ and $x_{b},~y_{b}$, satisfied $((x_{a}-x_{b})^{2}+(y_{a}-y_{b})^{2})^{1/2}<\Delta d$. Similarly, oppositely charged defects were identified as annihilation pairs if their final detection times and locations met the same criteria. The pair detection algorithm is designed to prevent more than one annihilation or creation event for any defect. Each defect’s trajectory is assigned a unique identifier, ensuring that the same identifier cannot be used in more than one pair. Trajectories that span fewer than five consecutive frames were excluded from the analysis. Events near the edge of the frame were excluded to prevent interference from off-frame defects. Additionally, defects appearing at the start or disappearing at the end of the time sequence were not considered created or annihilated to ensure accuracy. The detection algorithm identified $1679$ creation and $1676$ annihilation events in bacteria from three experimental replicas, and $1201$ creation and $1187$ annihilation events in HBECs from six experimental replicas.
Overall, the data on a single defect-pair is a sequence of vectors $\textbf{x}_{0},...,\textbf{x}_{K}$, where each vector $\textbf{x}_{i}$ is in $R^{8}$ ($x,y$ coordinates of each defect, one angle for $^{1}/_{2}$ and three angles for $^{1}/_{2}$). For a creation event, $\textbf{x}_{0}$ is the measurement closest to creation. For an annihilation event, $\textbf{x}_{K}$ is the measurement closest to annihilation. Furthermore, each pair-trajectory is classified as either “up” or “down”, based on the average projection of the $^{1}/_{2}$ direction on the direction connecting the pair (see sketch, Fig. 2). Specifically $\langle\varphi^{+}\rangle>0$ was defined as “up” and $\langle\varphi^{+}\rangle<0$ as “down”, where $\langle\cdot\rangle$ denotes the circular mean over the trajectory of the paired $^{1}/_{2}$ defect.
Trajectories of Creation and Annihilation
To describe the average trajectory of creation and annihilation events, consider the k’th trajectory with defined positions. The distance between defect-pairs is calculated as the Euclidean distance between the defect centers, $d=((x_{a}-x_{b})^{2}+(y_{a}-y_{b})^{2})^{1/2}$, where a and b refer to the to $^{1}/_{2}$ and $^{1}/_{2}$ defects, respectively. The angle of each defect, $\theta_{a/b}(t)$, was measured at various points along its trajectory relative to the position of the creation/annihilation event ($x_{0},y_{0}$). Since “up” and “down” defects rotate in opposite directions, we analyzed and averaged the trajectories of the “up” and “down” configurations separately. Finally, we combined the “up” and “down” trajectories by transforming the averaged “down” trajectory $\langle\theta_{a/b}(t)\rangle$ into $\langle-\theta_{a/b}(t)\rangle$.
Average Director and Flow Fields Around Defects
To measure the average fields around $\pm^{1}/_{2}$ defects, we analyzed all defect-pairs based on their configuration (“up” or “down”) and trajectory type (“creation” or “annihilation”), resulting in four distinct groups. Defects were analyzed alongside their corresponding flow fields within a window of $35~\mu m\times 35~\mu m$ for bacteria and $300~\mu m\times 300~\mu m$ for cells, centered at the defect core. The local director and velocity fields were then rotated to align all defects in the same orientation. Finally, the averaged, aligned fields were calculated for each defect type and group.
Quantifying irreversibility in single-defect trajectories
Here, we provide details for the statistical estimates presented in table 3.
For each $^{1}/_{2}$ defect, the detection algorithm described above provides a sequences of positions in 2D and orientation angle, $\theta\in[0,2\pi)$. The defect velocity $v$ is obtained from the displacements between consecutive frames (next frame - current frame). Denoting the velocity direction $\hat{V}=v/|v|$ and vectorial orientation $\hat{n}=(\cos\theta,\sin\theta)$, the projection of the defect orientation on the velocity direction (row 1 in table 3) is $\hat{V}\cdot\hat{n}$. The table shows the mean over all defects $\pm$ the standard error.
For each $^{1}/_{2}$ defect, the detection algorithm described above provides a sequences of positions in 2D and three orientation angles, corresponding to the directions of the three legs $\theta_{1},\theta_{2},\theta_{3}\in[0,2\pi)$. We define the average defect direction (taking into account the rank three tensor symmetry) as $p=\sum_{i}(cos\theta_{i},\sin\theta_{i})$. The angle corresponding to the defect orientation is then $\varphi=\tan^{-1}(p_{y}/p_{x})/3$. Note that a function such as atan2 should be used to obtain an angle in $0,2\pi)$. Finally, the projection of defect orientation in the velocity direction (angle $\psi$) is given by $\cos 3(\varphi-\psi)$. The table shows the mean over all defects $\pm$ the standard error.
[…]
See below for the estimation method for Kullaback-Leibler divergence.
Quantifying irreversibility in defect pairs
First, the data for each defect trajectory or defect-pair was reduced to include only a few relative angles, as described in the results section. Looking only at the first (last) $D$ samples right after (before) a creation (annihilation) event, each event corresponds to an effective vector $\textbf{f}_{i}$. For example, looking only at the relative angle between the $^{1}/_{2}$ and $^{1}/_{2}$ defects, each creation event is reduced to a $D$-dimensional vector $\textbf{f}_{i}=(\Theta_{1}\dots\Theta_{D})$, where $\Theta_{1}$ is the angle closest to creation. The time-reversed trajectory is simply $\textbf{b}_{i}=(\Theta_{D}\dots\Theta_{1})$.
The informatic entropy production is the Kullback-Leibler divergence between the distribution of forward observations $\textbf{f}_{i}$ and backward ones $\textbf{b}_{i}$. Hence, our experimental data provides us with samples from both spaces. In order to evaluate EP, we look at the sequence of projections and find the KLD between the density of forward and backward trajectories. In backward trajectories, the sequence of positions is inverted, and the velocity changes sign. Defect angles do not change; hence the projection also changes sign. Cutting trajectories into 2-frame snippets, we obtain a coarse-grained picture, which is a lower bound of the actual EP.
Finally, cross entropy between the annihilation and creation is defined as the sum of KLD(forward creation $||$ backward annihilation) + KLD(forward annihilation $||$ backward creation ). See below for the estimation method.
Estimation of Kullback-Leibler divergence
Let $F$ and $B$ denote continuous distributions over $\mathbb{R}^{D}$ such that $F$ is absolutely continuous with respect to $B$. The densities are denoted $f(x)$ and $b(x)$, respectively. We would like to estimate the Kullback-Leibler divergence of $F$ with respect to $B$,
$D_{\rm KL}(F||B)=\int_{\mathbb{R}^{D}}f(x)\ln\frac{f(x)}{b(x)}dx,$ (1)
using independent samples $\textbf{f}_{1}\dots\textbf{f}_{N}$ and $\textbf{b}_{1}\dots\textbf{b}_{M}$, taken from $F$ and $B$ respectively. Throughout, we take $0\ln 0/0=0$.
Here, the divergence is estimated using the $k$-nearest neighbors (kkNN) suggested by Wang et al Wang et al. (2009). The main idea is as follows. We note that the $D_{\rm KL}$ can be written as an average,
$D_{\rm KL}(F||B)=\left<\ln\frac{f(X)}{b(X)}\right$>, (2)
where $X$ denotes a random variable with distribution $F$. Let $k\ll N,M$. Locally, around a sample point $\textbf{f}_{i}$, the density $f(x)$ can be approximated as
$f(\textbf{f}_{i})\simeq\frac{k}{N-1}\frac{1}{V_{D}\rho_{k}^{D}(i)},$ (3)
where $V_{D}$ is the volume of the unit ball in $\mathbb{R}^{D}$ and $\rho_{k}(i)$ is the (Euclidean) distance from position $\textbf{f}_{i}$ of point $i$, to the $k$’s NN out of $\textbf{f}_{1}\dots\textbf{f}_{N}$ (not counting $i$ itself, hence N1N-1). Similarly, the density $b(x)$ around a sample point $i$, can be approximated as
$b(\textbf{f}_{i})\simeq\frac{k}{M}\frac{1}{V_{D}\nu_{k}^{D}(i)},$ (4)
$\nu_{k}(i)$ is the distance from $\textbf{f}_{i}$ to the $k$’s NN out of $\textbf{b}_{1}\dots\textbf{b}_{M}$. Substituting into (2) yields,
$D_{\rm KL}(F||B)\simeq\frac{D}{N}\sum_{i=1}^{N}\frac{\nu_{k}(i)}{\rho_{k}(i)}+\ln\frac{M}{N-1}.$ (5)
See Wang et al. (2009) for a proof (5) is a consistent estimator for $D_{\rm KL}(F||B)$.
To estimate the statistical error, we recall that the divergence is an average quantity. Accordingly (see also Beirlant et al. (1997) for similar considerations for estimation of entropy), we take the standard error to be
$\sigma_{\rm KL}(F||B)=\frac{1}{\sqrt{N}}{\rm std}\left[\ln\frac{f(X)}{b(X)}\right],$ (6)
where std denotes the standard deviation. Once again, approximating the densities using kkNN yield an estimator for the standard error Beirlant et al. (1997).
The parameter $k$, setting the number of neighbors, should be taken to be proportional to $\sqrt{N}$ Beirlant et al. (1997). Estimates reported for single-defects, single cells and the nematic/flow fields used $k=\sqrt{N}$. EP of defect pairs use only the orientation of defect, which are discretized. As a result, angles are not unique and the distence between neighbors may be zero. To this end, we use a larger value of k=2sqrtNk=2sqrt{N}.
Combining polar and nematic vector fields
The cells are long, thin objects with a shape that is broadly identical under $\pi$ rotation of the director. At high density, they form a thin layer with local orientational order. To describe the average orientation local of the we cells we introduce the director field $\hat{n}$ defined as:
$\underline{\hat{n}}=[\cos(\psi),\sin(\psi)]$ (7)
where $\psi$ is the average orientation of the cells at a point in space.
Most observations on the cells are compatible with the typical nematic active stress:
$\sigma_{ij}\sim(\hat{n}_{i}\hat{n}_{j}-\delta_{ij}/2)$ (8)
This is the lowest order active term consistent with conservation of momentum and nematic symmetry.
However, the cells here are able to self-propel, exchanging momentum with the substrate leading to an additional active body force on the material given. We introduce the second director field $\hat{p}$, which describes the self-propulsion direction of the cells. This can lead to a body force of the form:
$\underline{f}\sim\underline{\hat{p}}$ (9)
The cells are elongated in shape and self-propel along their long axis, meaning that $\underline{\hat{p}}$ and $\underline{\hat{n}}$ must be either parallel or anti-parallel, i.e. $(\underline{\hat{p}}.\underline{\hat{n}})^{2}=1$.
We introduce a polar director angle $\theta$ such that:
$\underline{\hat{p}}=[\cos(\theta),\sin(\theta)]$ (10)
Which implies that either $\theta=\psi$ or $\theta=\psi+\pi$.
Field around a topological defect
Topological defects within the nematic field are characterized by a half integer winding number $k=\pm^{1}/_{2}$. The director field at a polar angle $\phi$ around a topological defect is given by:
$\psi=k\phi+\psi_{0}$ (11)
where we have introduced a phase $\psi_{0}$ which corresponds to a global rotation of the defect by an angle of $\frac{\psi_{0}}{1-k}$. Note, here we will adopt the convention that $\phi\in[-\pi,\pi]$ for convenience and $\psi_{0}=0$ which fixes the orientation of the defect relative to the frame of reference. Examples of these defects are given in Extended Data Fig. 3.
A half integer winding number is incompatible with the polar field $\hat{p}$, and necessitates the introduction of a discontinuous jump in the polar director angle, $\theta$. This is the grain boundary over which $\underline{\hat{p}}$ goes through a $\pi$ rotation which begins and ends at a topological defect. We introduce the grain boundary as follows
$\theta=k\phi+\psi_{0}+\pi*(H(\phi-\phi_{g})+m)$ (12)
$m\in\{0,1\}$ corresponds to a global rotation of $\hat{p}$ by $\pi$ and captures the additional symmetry of the polar field relative to a nematic field. The grain boundary is introduced with a Heaviside step function given by:
[…]
$\phi_{g}$ describes the polar angle at which the grain boundary approaches the core of the topological defect. Without loss of generality, we constrain $\phi_{g}\in(-\pi,\pi]$. The grain boundary and resulting polar field are shown in Extended Data Fig. 3.
Total force on a topological defect core due to polar forcing
We now calculate the total additional force on the defect core due to polar forcing by integrating the polar force azimuthally around the core of the defect. Without loss of generality we set $\psi_{0}=0$, which fixes the orientation of the defect. This results in the tail of the $^{1}/_{2}$ defect, or one of the legs of the $^{1}/_{2}$ defect, pointing in the $\phi=0$ direction, which we set as the $\hat{\underline{x}}$ direction.
The grain boundary breaks the mirror symmetry of a $^{1}/_{2}$ defect if is introduced at any angle other than that opposite to the tail of the defect. Thus in most cases, the $^{1}/_{2}$ defect becomes chiral.
The total body force around the core of the defect with charge $k$ is given by.
$\underline{F}$ $=\oint\underline{f}(\phi)\textrm{d}\phi$ (14)
$\sim\oint\underline{\hat{p}}(\phi)\textrm{d}\phi$ (15)
It is convenient to evaluate this integral piecewise to find the solution:
$\underline{F}$ $\sim\int_{-\pi}^{\phi_{g}}\underline{\hat{p}}(\phi)\textrm{d}\phi+\int_{\phi_{g}}^{\pi}\underline{\hat{p}}(\phi)\textrm{d}\phi$
$=\int_{-\pi}^{\phi_{g}}[\cos(k\phi+m\pi),\sin(k\phi+m\pi)]\textrm{d}\phi+$
$+\int_{\phi_{g}}^{\pi}[\cos(k\phi+(m+1)\pi),\sin(k\phi+(m+1)\pi)]\textrm{d}\phi$
$=\left[\frac{[\sin(k\phi+m\pi),-\cos(k\phi+m\pi)]}{k}\right]_{-\pi}^{\phi_{g}}+$
$+\left[\frac{[\sin(k\phi+(m+1)\pi),-\cos(k\phi+(m+1)\pi)]}{k}\right]_{\phi_{g}}^{\pi}$
$=\frac{2}{k}\left[\sin\left(k\phi_{g}+m\pi\right),-\cos\left(k\phi_{g}+m\pi\right)\right]$
$=\frac{2}{k}\left[\cos\left(k\phi_{g}+(m-0.5)\pi\right),\sin\left(k\phi_{g}+(m-0.5)\pi\right)\right]$
This force is shown as the large red arrow in Extended Data Fig. 3. Some important things to note here.
First, there is a net translational force on both $\pm^{1}/_{2}$ defects, meaning that we expect both defects to self-propel. This is in addition to the self-propulsion force already present for $^{1}/_{2}$ defects due to the nematic active stress.
If $\phi_{g}=0$, as is the case immediately after creation, the additional force is perfectly perpendicular to the symmetry axis of the defect. This would imply a self-propulsion perpendicular to the tail for $^{1}/_{2}$ defects (or a leg for $^{1}/_{2}$ defects).
The additional force on the defects breaks chiral symmetry provided it does not align with a mirror symmetry axis of the defect. The mirror symmetry axes of the defect are given by $\phi_{s}=n\pi/2(k-1)$, $n\in\mathbb{Z}$. In addition, $m$ rotates the force by $\pi$ meaning it changes the sign of the additional force. This switches the chirality of the effect. As mentioned in the main text, $m$ arises from a spontaneously broken symmetry.
Total torque on a topological defect core due to polar forcing
We can assess the active torsional force around a defect caused by the polar active force by azimuthally integrating the azimuthal component of the polar force. This is given by:
$\tau=\oint\underline{f}.\underline{\textrm{d}l}$ (17)
This integration is performed on a circle around the defect and $\underline{dl}$ is tangent to that circle. Thus we can write:
$\tau\sim\int_{-\pi}^{\pi}\underline{\hat{p}}.[-\sin(\phi),\cos(\phi)]\textrm{d}\phi$ (18)
We can simplify this with angle addition formula to obtain:
$\tau$ $\sim\int_{-\pi}^{\pi}\sin(\theta-\phi)\textrm{d}\phi$ (19)
$=\int_{-\pi}^{\pi}\sin((k-1)\phi+\pi*(H(\phi-\phi_{g})+m))\textrm{d}\phi$ (20)
$=\left[\frac{-\cos((k-1)\phi+m\pi)}{k-1}\right]_{-\pi}^{\phi_{g}}+$ (21)
$+\left[\frac{-\cos((k-1)\phi+(m+1)\pi)}{k-1}\right]_{\phi_{g}}^{\pi}$ (22)
$=\frac{1}{k-1}\left[2\cos((k-1)\phi_{g}+(m+1)\pi)+\right$. (23)
$\quad\left.\cos(m\pi)-\cos((2k+m-1)\pi)\right]$ (24)
$=\frac{2}{1-k}\cos((k-1)\phi_{g}+m\pi)$ (25)
This torque is shown as the large green arrow in Extended Data Fig. 3.
Some things to note here:
First, the sign of the torque depends on $m$, thus is selected by a spontaneous symmetry break.
Second, the magnitude of the torque depends on the orientation of the grain boundary. The torque disappears when $(k-1)\phi_{g}+m\pi=(n+0.5)\pi$, $n\in\mathbb{Z}$. This is equivalent to when the net additional net force aligns with an axis of miror symmetry.
Third, the magnitude of the torque additionally depends on $k$, with $^{1}/_{2}$ defects experiencing three times more torque on their core than $^{1}/2$ defects.
IV Supplementary material
Table 3: Entropy production (EP) estimation - control. Top: EP for trajectories of individual bacteria and cells. Bottom: EP for the nematic and flow fields at random positions. Values show the estimate $\pm$ the standard error. Since this is a lower bound, a value that is statistically larger than zero implies non reversible dynamics. See the methods section for details.
Bacteria Cells
EP individual trajectories $004\pm 0.004 0.0$$003\pm 0.0002$
EP nematic+flow $008\pm 0.001 0.000$$008\pm 0.001$

### v2

@@
 However, in contrast to active nematic theory, we find that defect creation and annihilation undergoes spatial symmetry breaking.
-We propose that this results from a fundamental dualism between nematic structural organization and generated polar forces, which are intrinsic to living systems.
-Furthermore, our estimation of entropy production reveals that creation and annihilation are not reversed processes.
+We propose that this results stems from a fundamental dualism between nematic structural organization and generated polar forces, which are intrinsic to living systems.
+Furthermore, estimation of entropy production reveals that creation and annihilation are not reversed processes.
 Our findings challenge conventional nematic models and emphasize the role of defect-mediated dynamics in non-equilibrium biological systems as a major source of entropy production.
@@
 Elasticity arises from an elongated cellular shape, leading to liquid-crystal-like properties Be’er and Ariel (2019); Aranson (2022); Doostmohammadi and Ladoux (2022).
-The resulting defects are mesoscopic objects involving tens to hundreds of cells and governed by intrinsic spatiotemporal scales that are independent of system size Li et al.
+The resulting defects are meso-scopic objects involving tens to hundreds of cells and governed by intrinsic spatiotemporal scales that are independent of system size Li et al.
 (2019); Yashunsky et al.
@@
 (2023).
-Recent studies suggested that nematic symmetry can also manifest in systems with directionality Venkatesh et al.
-(2025), such as models of self-propelled rods Shi and Chaté (2018); Großmann et al.
-(2020) and bacterial experiments Head et al.
+Recent studies suggested that nematic symmetry can also manifest in systems with clear polar directionality Venkatesh et al.
+(2025), such as self-propelled rod models Shi and Chaté (2018); Großmann et al.
+(2020); de Graaf Sousa et al.
+(2025) and rod-like bacteria gliding on surfaces Han et al.
 (2025).
-Intriguingly, some nematically ordered systems display polar characteristics under specific conditions, indicating the coexistence of nematic and polar symmetries Huber et al.
+Under certain conditions, active nematic systems can even acquire polar features, indicating that nematic and polar symmetries can coexist Huber et al.
 (2018); Meacock et al.
@@
 (2022); Lacroix et al.
-(2024).
-This interplay of symmetries is fundamental to understanding biological processes, such as biofilm formation and the organization of cellular tissues, yet remains underexplored experimentally.
+(2024); Vafa and Doostmohammadi (2025).
+In Han et al.
+(2025), polar stresses in the nematically organized Myxococcus xanthus colonies were shown to be associated with multi-layer formation.
+Understanding how living systems organize may depend on this interplay of symmetries, which could underlie processes such as biofilm formation and tissue organization, yet it remains largely untested experimentally.
 Defect interactions have been primarily investigated through numerical simulations Thampi et al.
@@
 (2021).
-Elastic forces between defects predict Coulomb-like interactions, as well as additional forces and torques that rely on their relative orientation Shankar et al.
+Elastic forces between defects result in polar-like interactions, as well as additional forces and torques that rely on their relative orientation Shankar et al.
 (2022); Pearce and Kruse (2021); Tang and Selinger (2017); Vromans and Giomi (2016).
-It has been suggested that nematic active stresses also lead to the generation of hydrodynamic flows around the defects which results in self-propulsion of $^{1}/_{2}$ defects.
+It has been observed that nematic active stresses also lead to the generation of hydrodynamic flows around the defects which results in self-propulsion of $^{1}/_{2}$ defects.
 Rotational diffusion can reduce persistence in $^{1}/_{2}$ defect motion, preventing unbinding from $^{1}/_{2}$ counterparts, and affecting defect density in steady states Shankar et al.
@@
 Recent experiments on microtubule-based active nematics and swarming bacteria have uncovered a broad spectrum of defect organization, ranging from giant number fluctuations to hyperuniformity Yashunsky et al.
-(2024); Fernandez-Nieves et al.
+(2024); de la Cotte et al.
 (2025); Pearce et al.
@@
 (2024).
-However, the creation and annihilation processes, which we identified below as primary sources of energy injection and dissipation in living systems, are underexplored and necessitating experimental validation of current theoretical predictions.
+However, the creation and annihilation processes, which we identify below as primary sources of energy injection and dissipation in living systems, are underexplored, necessitating experimental validation of current theoretical predictions.
 In this paper, we experimentally study swarming bacteria Be’er and Ariel (2019) and human bronchial epithelial cells (HBECs) Blanch-Mercader et al.
-(2018)—distinct systems that both display nematic half-integer defects.
+(2018)—two distinct systems that are known to display nematic half-integer defects.
 Our findings reveal that the creation and annihilation of defects follow distinct, previously unrecognized trajectories that spontaneously break mirror symmetry—a feature absent in standard active nematic models.
 We also find that these trajectories exhibit time-reversal symmetry breaking, a hallmark of non-equilibrium processes.
-Statistical analysis reveals that defect creation and annihilation drive substantial entropy production, underscoring their irreversible nature and critical role in active matter organization.
-Finally, we demonstrate why these observations can be explained by taking into account directional (polar) self-propulsion in bacteria and cells.
+Statistical analysis reveals that defect creation and annihilation are a major source of entropy production in the bulk, underscoring their irreversible nature and critical role in active matter organization.
+Finally, we propose a theoretical model that explains these observations taking into account directional (polar) self-propulsion in bacteria and cells.
+II Results
 Figure 1: Director field and nematic defects in swarming Bacillus subtilis and human bronchial epithelial cells (HBECs) monolayers.
@@
 subtilis and (d) HBECs.
-II Results
 Half-integer defects in bacteria and cells
@@
 Despite the extreme disparity in the time scales and dimension (Fig.
-1), both systems share fundamental characteristics of active extensile nematics Giomi (2015); Mart´ınez-Prat et al.
+1), both systems share fundamental characteristics of active extensile nematics Giomi (2015); Martínez-Prat et al.
 (2021): (i) A steady-state mixture of half-integer defects (Figs.
@@
 1c-d), implying an approximately constant number of defects over time with a total charge that fluctuates around zero.
-We characterize the motion of defect pairs through these observables, see the sketch in Fig.
-2.
+We characterize the motion of defect pairs through these observables, see the sketches in Figs.
+2 and 3.
   • $t$: time from creation or annihilation event (negative time indicates time before annihilation).
@@
   • $\varphi^{+}$: the angle between the orientation of the “tail” of the comet-shaped $^{1}/_{2}$ defect and the vector connecting the pair.
-Thus, $\varphi^{+}=0$ indicates that the tail is pointing away from the $^{1}/_{2}$ defect.
-  • $\varphi^{-}$: the angle between the orientation of one of the $^{1}/_{2}$ legs and the line connecting the pair.
+Thus, $\varphi^{+}=0$ indicates that the tail is pointing toward the $^{1}/_{2}$ defect.
+  • $\varphi^{-}$: the angle between the orientation of one of the $^{1}/_{2}$ legs and the vector connecting the pair.
 We use the smallest angle out of the three legs.
-For example, $\varphi^{-}=0$ implies that one of the legs is pointing towards the $^{1}/_{2}$ defect.
+For example, $\varphi^{-}=0$ implies that one of the legs is pointing away from the $^{1}/_{2}$ defect.
   • $\delta=\varphi^{+}-3\varphi^{-}-\pi$: pair orientation offset.
 This gives a measure of how compatible the orientations of the defects are with each other.
-$\delta=0$ suggests the defects are at the orientations that minimize the variation in the nematic field between themPearce and Kruse (2021).
+$\delta=0$ suggests the defects are at the orientations that minimize the variation in the nematic field between them Pearce and Kruse (2021); Tang and Selinger (2017); Vromans and Giomi (2016).
 Figure 2: Configuration of $^{1}/_{2}$ and $^{1}/_{2}$ defect-pairs during creation and annihilation.
-Sketch: Definitions of orientational angles $\varphi^{+}$ (pink) and $\varphi^{-}$ (violet) for paired $^{1}/_{2}$ and $^{1}/_{2}$ defects relative to the pair axis connecting between $^{1}/_{2}$ and $^{1}/_{2}$ defect centers (horizontal black line).
+Center sketch: Definitions of orientational angles $\varphi^{+}$ (pink) and $\varphi^{-}$ (violet) for paired $^{1}/_{2}$ and $^{1}/_{2}$ defects.
+These angles are defined relative to the vector pointing from the $^{1}/_{2}$ to the $^{1}/_{2}$ defect, shown by the solid black arrow.
 (a, b) Distribution of the pair orientation offset angle, $\delta=\varphi^{+}-3\varphi^{-}-\pi$, as a function of time for creation (blue) and annihilation (red) events.
-$\delta=0$ implies that the nematic director fields of the paired-defects are synchronized.
+$\delta=0$ implies that the nematic director fields of the paired-defects are in phase.
 (c, d) Distribution of $\varphi^{+}$ as a function of time for the creation (blue) and annihilation (red) events.
@@
 Red and blue color intensities indicate probability density.
-(e-g) Alignment of defect pairs, classified by the mean orientation of the $^{1}/_{2}$ defect during its trajectory, $\langle\varphi^{+}\rangle$.
-Angular distributions of $\varphi^{+}(d<r_{0})$ and $\varphi^{-}(d<r_{0})$ within $r_{0}=15~\mu m$ for bacteria and $35~\mu m$ for cells are shown for “up” ($\langle\varphi^{+}\rangle>0$, green) and “down” ($\langle\varphi^{+}\rangle<0$, yellow) configurations in bacteria (e, f) and cells (g, h).
-Spontaneous breaking of the mirror symmetry in defect orientation
-At annihilation or creation events, defect pairs are in phase ($\delta=0$; see Figs.
+The bright line close to $t=0$ indicates a region in which the distance between the defects is up to half cell-body length ($3-4~\mu m$ for bacteria and $7-8~\mu m$ for cells).
+At these distances the continuous nematic field is not well defined and the discrete nature of particles is of concern.
+(e-g) Angular distribution of $\varphi^{+}$ and $\varphi^{-}$ for defect pairs at small separation ($d<15~\mu m$ for bacteria and $d<35~\mu m$ for cells).
+Pairs are classified as “up” ($\langle\varphi^{+}\rangle>0$, green) or “down” ($\langle\varphi^{+}\rangle<0$, yellow).
+Spontaneous breaking of mirror symmetry during defect creation and annihilation
+At the point of annihilation or creation, defect pairs are in phase ($\delta=0$; see Figs.
 2a-b), indicating they are perfectly aligned.
-When $\delta~\neq 0$, the elastic energy between the defects diverges at small separations, indicating that the defects align at very short distances Pearce and Kruse (2021).
-We next look at the alignment of the $^{1}/_{2}$ defect with the line connecting the defects in the pair, captured by the parameter $\varphi^{+}$.
-We see a bimodal distribution for $\varphi^{+}$ immediately before annihilation with peaks at $\pm\pi/2$; see Figs.
+This is due to elastic torques that diverge when defects are out of phase at small distances Pearce and Kruse (2021).
+However, the orientation of the $^{1}/_{2}$ defect relative to the position of the $^{1}/_{2}$ defect (given by $\varphi^{+}$) is free to vary, provided they remain in phase.
+In a typical active nematic system, the average orientation of the $^{1}/_{2}$ defect during creation or annihilation is dictated by the active stress, which causes the $^{1}/_{2}$ defect to self-propel; for an extensile active nematic one expects $\varphi^{+}=0$ during creation and $\varphi^{+}=\pm\pi$ during annihilation Giomi et al.
+(2014); Shankar et al.
+(2018).
+Conversely, in both our bacteria and cellular experiments, we observe a bimodal distribution of $\varphi^{+}$ immediately before annihilation with peaks around $\pm\pi/2$; see Figs.
 2c-d.
-This indicates that the tail of the positive defect is perpendicular to the direction of annihilation at the end of this process.
+This indicates that the tail of the positive defect is oriented roughly perpendicular relative to the direction of annihilation at the end of this process.
+Since $\delta=0$, $\varphi^{-}$ varies by about one-third as much.
 A similar but less pronounced effect is also visible immediately after creation.
-This bimodality indicates a spontaneously broken chiral symmetry in the orientation of the defects.
-To quantify this effect, we split all trajectories into two groups based on their time-averaged orientation across the full trajectory, either “up” (average $\varphi^{+}>0$) or “down” (average $\varphi^{+}<0$).
+This bimodality indicates a spontaneously broken mirror symmetry in the process of creation and annihilation.
+To quantify this effect, we split all trajectories into two groups based on their time-averaged orientation across the full trajectory, either “up” ($\langle\varphi^{+}\rangle>0$) or “down” ($\langle\varphi^{+}\rangle<0$).
 The orientation $\varphi^{-}$ follows a similar symmetry breaking, consistent with $\delta=0$ (Figs.
-2e-f).
+2e-g).
+In contrast, active tubulin-kinesin suspensions, which are purely nematic, exhibit no such mirror-symmetry breaking behavior.
+Analysis of data from Tan et al.
+Tan et al.
+(2019) (Extended Figs.
+4c and 5) shows that the distribution of $\varphi^{+}$ has a single peak around 0 (creation) or $\pi$ (annihilation), in accordance with the prediction of active nematic theory.
 Figure 3: Spiraling trajectories of creation and annihilation.
-Sketch: Co-rotation angle $\gamma$—the line connecting between the location of defect centers with respect to the angle close to creation/annihilation instant ($\gamma_{d\rightarrow 0}$).
+Center sketch: Co-rotation angle $\gamma$—the line connecting between the location of defect centers with respect to the angle close to the creation/annihilation instant ($\gamma_{d\rightarrow 0}$).
 (a, b) Distance between $\pm^{1}/_{2}$ defect-pairs as a function of time after creation (blue) and before annihilation (red), for bacteria and cells.
@@
 (e, f) Average trajectories of $^{1}/_{2}$ (pink) and $^{1}/_{2}$ (violet) defects during creation and annihilation, measured in the lab frame of reference.
-Trajectories are centered around the creation/annihilation position (black point), rotated to $\gamma_{d\rightarrow 0}$ and mirrored to the “up” configuration.
+Trajectories are centered around the creation/annihilation position (black point), rotated to $\gamma_{d\rightarrow 0}=0$ and mirrored to the “down” configuration.
+Figure 4: Symmetry breaking in the flow around $^{1}/_{2}$.
+(a–b) The distribution of $\phi_{V}$ in bacteria, the angle between the $^{1}/_{2}$ velocity and defect orientation (inset in panel a).
+Average flow fields around $^{1}/_{2}$ defects are obtained by separately averaging trajectories according to their defect-pair configurations: “up” ($\langle\varphi^{+}\rangle>0$) or “down” ($\langle\varphi^{+}\rangle<0$).
+Flows are shown for defects with overall displacement to the right (blue–gray) and to the left (pink).
+The velocity scale indicated by the black arrow is $10~\mu\mathrm{m}/\mathrm{s}$.
+Scale bar: $10~\mu\mathrm{m}$.
 Spiraling trajectories of creation and annihilation
 The separation rate between paired defects (Figs.
-3a-b) is characterized by increased speed close to the creation or annihilation events, consistent with theoretical predictions for defect annihilation in passive systems Pearce and Kruse (2021).
-However, defects do not move in a straight line.
-This is quantified by the orientation of the vector connecting the defects changes over time, $\gamma(t)$.
+3a-b) is characterized by increased speed close to the creation or annihilation events, consistent with theoretical predictions for defect annihilation Giomi et al.
+(2014).
+However, contrary to typical active nematics, defects do not move in a straight line, but rather spiral toward or away from each other, see Figs.
+3e-f for average trajectories of defects in the “down” configuration.
+The handedness of the spiral motion can be quantified by plotting $\gamma(t)$; the orientation of the vector connecting the defects.
 Generally, $\gamma$ varies monotonically with time, indicating that defect pairs tend to co-rotate during the process of annihilation or creation.
-Furthermore, the chirality of the co-rotation is correlated with the “up” or “down” configuration (Figs.
+Furthermore, the direction of the co-rotation is correlated with the “up” or “down” configuration (Figs.
 3c-d).
-Lastly, Figs.
-3e-f show averaged trajectories for creation and annihilation events in the “up” configuration.
-This reflects the spiraling motion toward (away) each other during annihilation (creation).
-These trajectories are similar to those predicted for the annihilation of topological defects with an initial phase disparity Pearce and Kruse (2021).
-Figure 4: Ensemble-averaged flow fields near $\pm^{1}/_{2}$ defect pairs.
-(a, b) Average velocity fields (black arrows) and vorticity distributions (heatmap) around defects.
-(a-d) Flow fields around paired $^{1}/_{2}$ and $^{1}/_{2}$ defects, averaged separately for the “up” (a, b) and “down” (a, d) configurations.
-e, f) Flow fields around unpaired $^{1}/_{2}$ and $^{1}/_{2}$ defects.
-For clarity, only results for bacteria close to annihilation are shown.
-The angular orientation of defects is illustrated in the schematic.
-Arrow and spatial scales are indicated.
-Scale bar: $10~\mu m$.
-The number of fields averaged to obtain each panel is: (a) $5459$; (b) $5391$; (c) $5336$; (d) $5323$; (e) $89027$; (f) $88267$.
-Spontaneous breaking of the mirror symmetry in flow fields
-We find that flows around defect pairs also exhibit mirror symmetry breaking (Figs.
-4a-d, Extended Data Figs.
-1-2).
-Flows were measured separately for the “up” and “down” orientations.
-In bacteria, the forward flow at the core of the defect $^{1}/_{2}$ is offset relative to the defect symmetry axis.
-The direction of the tilt changes between annihilation and creation (Extended Data Fig.
-1).
-In particular, for both bacteria and cells, the flow at the core of the $^{1}/_{2}$ defects exhibits a distinct rotational behavior, rotating clockwise for “up”-oriented configurations (Fig.
-4B, Extended Data Figs.
-1b,f and 2b,f) and counterclockwise for “down”-oriented configurations (Fig.
-4d, Extended Data Figs.
-1d,h and 2d,h).
-The flows around unpaired defects (Figs.
-4e-f) retain mirror symmetry, in agreement with theoretically predicted flows around isolated defects in extensile active nematics.
+These spiral paths are similar to those predicted for the annihilation of topological defects with an initial phase disparity Pearce and Kruse (2021); Vromans and Giomi (2016), however since we do not observe any major phase disparity (Figs.
+2a-b) we predict that the spiraling arises from active forces and self-propulsion of the $^{1}/_{2}$ defect.
+Spontaneous breaking of mirror symmetry in the flow fields
+self-propulsion of $^{1}/_{2}$ defects arises from the characteristic flow fields generated around the defect by the active forces.
+For a typical active nematic, the flow around a $^{1}/_{2}$ defect features two counter-rotating vortices and a central flux aligned with the defect which causes it to self-propel Giomi et al.
+(2014).
+In order to quantify the direction in which defect move we introduce two additional angles to be measured.
+  • $\phi_{V}$: For a single $^{1}/_{2}$ defect, $\phi_{V}$ is the angle between the defect velocity (calculated from the displacement between consecutive frames) and the defect orientation.
+In words, $\phi_{V}$ is the offset between the orientation of a $^{1}/_{2}$ defect and the direction in which it moves.
+The theory of active nematics predicts that for an isolated defect $\phi_{V}=\pm\pi$ Giomi et al.
+(2014).
+Figure 4 shows the direction of the $^{1}/_{2}$ movement, $\phi_{V}$, during creation and annihilation events.
+The distribution of angles is bimodal, suggesting that the spontaneous break in symmetry is also manifested in a slanted trajectory (compared to the defect orientation).
+This can be directly observed by averaging the flows around recently created/soon to be annihilated $^{1}/_{2}$’s, partitioned according to their “up” or “down” configurations (Fig.
+4).
+We find low, yet highly significant correlation between $\phi_{V}$ close to creation and close to annihilation of the same defect $^{1}/_{2}$ (for details see Extended Data Fig.
+6).
+This indicates that the direction off-axis motion is consistent throughout the lifetime of the defect.
+Overall, the results above suggest that $^{1}/_{2}$ are created with an “up” or “down” orientation.
+This orientation persists throughout the life of the defect and finally determines its configuration during annihilation.
+The configuration of the $^{1}/_{2}$ seems to be constrained or enslaved to the $^{1}/_{2}$, Extended Data Figs.
+1 and 2
+Finally, comparing with active nematic tubulin-kinesin suspensions, which do not exhibit mirror-symmetry breaking, the displacement of $^{1}/_{2}$ defects remains aligned with their orientation.
+This is evidenced by the single peak in the distribution of $\phi_{V}$ at $\pi$ (Extended Fig.
+6).
+This demonstrates that the mirror-symmetry breaking observed in bacteria and cells does not arise from the presence of counter-rotating vortex structures, which are generic to active nematic turbulence Doostmohammadi et al.
+(2016).
 Figure 5: The interaction between nematic and polar symmetries.
-(a) Defect creation via bend instability with the overlaid polar field, indicated by red arrows; a grain boundary in the polar field is introduced indicated by the blue dashed line.
-The polar field leads to a torque on the defects, indicated by the green arrow.
-The torque will reorient the $^{1}/_{2}$ defect, causing it to self-propel around the $^{1}/_{2}$ defect.
-Shown here is the “up” configuration.
+(a) Defect creation via bend instability with the overlaid nematic (black lines) and polar (red arrows) fields.
+A grain boundary in the polar field is introduced (blue dashed line).
+The polar field leads to a torque on the defects (green arrow), which will reorient the $^{1}/_{2}$ defect, causing it to self-propel and spiral around the $^{1}/_{2}$ defect (inset).
+Shown here is the “down” configuration.
+$\zeta_{p}\neq 0$.
 (b) Nematic and polar fields around a pair of defects before annihilation.
-Here the grain boundaries do not directly connect.
-The same torque is present around the $^{1}/_{2}$ defect that can lead to a spiral trajectory.
-(c) & (d) Predicted Stokes flow field around a (c) $^{1}/_{2}$ and (d) $^{1}/_{2}$ defect with additional active force along the polar direction; shown in the “up” configuration.
+Here, the grain boundaries do not directly connect.
+A similar torque, acting on the $^{1}/_{2}$ defect, leads to a spiraling trajectory (inset).
+(c) Total co-rotation ($\Delta\gamma$) and (d) final positive defect orientation ($\Delta\varphi^{+}$) of the defect pair as a function of the applied polar body force ($\zeta_{p}$).
+Dashed lines indicate creation and solid lines for annihilation, green “up” and yellow “down” configuration.
+(e) Predicted Stokes flow around a $^{1}/_{2}$ defect in the annihilation “up” (or creation “down”) configuration.
+Shown with $|\zeta_{p}/\zeta_{Q}|=0.5$ and $\phi_{g}=0$.
+(f) Angular offset of defect self-propulsion ($\phi_{V}$) and (g) net vorticity around a $^{1}/_{2}$ defect as a function of the relative polar body force ($\zeta_{p}/\zeta_{Q}$) and position of the grain boundary ($\phi_{g}$).
+$\phi_{g}=0$ implies that the grain boundary alignment with the tail of the $^{1}/_{2}$ defect is as in (a) and (b).
 The source of mirror symmetry breaking
-In addition to nematic order, bacteria and cells also exhibit a polar symmetry due to their directed self-propulsion mechanism.
-Combining nematic interactions with active stresses of various symmetries can significantly alter active nematic behavior Ardaševa et al.
-(2025); Großmann et al.
-(2020).
 To describe cell self-propulsion, we introduce a polar field, $\underline{p}$, indicating the direction of cell movement relative to the substrate, alongside the nematic field that describes the cells’ long axes.
@@
 2c-d.
-Additionally, the reverse of this process is equally likely, resulting in the observed “up” and “down” configurations during annihilation.
+Additionally, the reverse of this process is equally likely, resulting in the observed “up’ and “down’ configurations during annihilation.
+In this section we provide a possible physical explanation to the symmetry breaking observed above.
+Cellular systems such as those studied here are often analyzed using the framework of active nematics with good agreement.
+In particular, the constant turnover of half integer defects, in which $^{1}/_{2}$ defects are created through bend instability and self-propulsion, indicates that the system behaves as an extensile active nematic.
+However, in addition to nematic order, the bacteria and cells studied here also exhibit polar symmetry through their directed self-propulsion mechanism Zhou et al.
+(2014); Be’er and Ariel (2019); Vaidžiulytė et al.
+(2022); Lo Vecchio et al.
+(2024); Lacroix et al.
+(2024).
+Combining nematic interactions with active stresses of various symmetries can significantly alter active nematic behavior Patelli et al.
+(2019); Ardaševa et al.
+(2025); Großmann et al.
+(2020); de Graaf Sousa et al.
+(2025).
+In particular, breaking the polar symmetry of the nematic field at the microscopic level can lead to a break in chiral symmetry at the macroscopic level.
+To describe the orientation of the cells we utilize a nematic tensor field, $\underline{\underline{Q}}$, which governs the steric interactions between the cells.
+In addition, we describe the head-tail symmetry of the cells by a polar vector field, $\underline{p}$, which governs the self-propulsion of the cells against the substrate.
+This can be visualized as two superimposed parallel fields: one with lines describing the long axes of the cells, and another with arrows for the head direction (see Fig.
+5a).
+Both $\underline{p}$ and $\underline{p}$ are compatible with the same nematic field.
+We model the system using an active nemato-polar framework in which we simultaneously simulate $\underline{\underline{Q}}$ and $\underline{p}$.
+The dynamics of $\underline{p}$ are constrained to follow the dynamics of $\underline{\underline{Q}}$ keeping the two fields parallel.
+To generate the motion of the cells we include two active terms: a standard extensile active stress, $\sigma^{a}=\zeta_{Q}\underline{\underline{Q}}$, and a body force describing cellular self-propulsion against the substrate, $\underline{f}=\zeta_{p}\underline{p}$.
+See SI for details.
+In a typical extensile active nematic, defects are created when a bend distortion in the director field becomes unstable.
+The bend distortions are characterized by a non-zero curl, which is unsigned for the nematic field.
+However, in the polar field the bend distortion is signed and therefore breaks mirror symmetry.
+The resulting half integer defects are incompatible with the polar field.
+This leads to a grain boundary that connects the two defects, across which the polar field changes sign, see blue line on Fig.
+5a.
+Notably, the mirror image of this process produces defects with opposite chirality and is equally likely, indicating that defect chirality selection is driven by spontaneous symmetry breaking during creation.
+In this context, distinct “up” and “down” defect configurations can emerge during creation.
+We simulate the defect creation process by initializing the fields with a single prominent bend distortion at the center of the simulation domain.
+Upon simulation, the bend distortion nucleates a pair of defects that separate and follow curved trajectories, see Fig.
+5a (inset), recreating those observed in experiments, see Fig.
+3e-f.
+We simulate the annihilation process by initializing the simulation with a pair of defects such that $\varphi^{+}=\pm\pi$, (see Fig.
+5b), as expected in active nematic.
+Simulated annihilations also show curved trajectories (Fig.
+5b inset), resembling the experimental observations (Fig.
+3e-f).
+We quantify the co-rotation of the defect pair by plotting the total change in $\gamma$ throughout the creation or annihilation process, see Fig.
+5c.
+We see that the degree of co-rotation increases with the strength of the polar forcing, $\zeta_{p}$, and is larger during annihilation.
+We also observe rotation of the $^{1}/_{2}$ defect, quantified by $\Delta\varphi^{+}$ at the end of the simulation, shown in Fig.
+5d.
+The rotation of the defect increases with the magnitude of the polar force.
+Consistent with our experimental observations (Fig.
+3e–f), the rotation is more pronounced during annihilation than during creation.
+Importantly, the chirality of the simulations can be reversed either by mirroring the fields or, equivalently, by changing $\zeta_{p}\rightarrow-\zeta_{p}$.
+This procedure allows to simulate the system in either the “up” or “down” configuration, shown as green or yellow in Figs.
+5e-f.
+If we remove the active body force, $\zeta_{p}=0$, we recover the straight trajectories expected for typical active nematics.
+In our experiments, the flow profile around a $^{1}/_{2}$ defect is offset relative to the symmetry axis of the defect, see Fig.
+4.
+We can analytically calculate how the introduction of the polar body force influences the flow around a $^{1}/_{2}$ defect and observe a similar offset, see Fig.
+5e and SI for details.
+The polar body force breaks the mirror symmetry of the defect, resulting in a central flux at an angle $\phi_{V}$ relative to the tail of the $^{1}/_{2}$ defect.
+The magnitude of this angular offset depends on the angular position of the grain boundary ($\phi_{g}$) and the scale of the polar active force, see Fig.
+5f.
+The distribution of $\phi_{V}$ measured from our experiments has a peak at $\approx\pm\pi/4$ for bacteria and $\approx\pm\pi/3$ for cells, see Fig.
+4 and Extended Fig.
+6a-b.
+Such an offset would indicate that the polar forcing accounts for at least $25\%$ of the flow around the defect.
+The polar body force also introduces a net vorticity around the core of the defect, see Fig.
+5g.
+While the net flux at the core of the defect causes it to self-propel, the net vorticity induces a torque that causes defects to rotate.
+Similar to the velocity offset, the magnitude of the rotation is controlled by the magnitude of the polar active force, but the direction of rotation depends only on the sign of $\zeta_{p}$, hence the chirality of the defect.
+Taken together, these observations explain the spontaneous symmetry breaking into “up” and “down” configurations observed in experiments.
+Defects are generated through a bend instability due to the extensile active stress, hence with $\varphi^{+}\approx 0$.
+This process breaks the mirror symmetry of the polar field and introduces a grain boundary across which $\underline{p}$ changes sign (blue dashed line in Fig.
+5a).
+The grain boundary connects the defects and results in a net curl around the $^{1}/_{2}$ defect which translates to a net torque (green arrow in Fig.
+5a).
+This torque causes the $^{1}/_{2}$ defect to rotate as it self-propels, thus $\dot{\varphi}^{+}\neq 0$.
+Combined with the defect self-propulsion, this rotation results in the defect taking a curved trajectory, spiraling away from its $^{1}/_{2}$ equivalent (Fig.
+5a).
+The sign of the torque depends on the sign of the curl of the initial bend distortion.
+Thus, it arises from a spontaneous break in symmetry.
+A counter clockwise net torque on the defect leads to $\dot{\varphi}^{+}>0$ and we observe the “up” configuration, as shown in Fig.
+5a.
+The polar grain boundary connected to the defect core is topologically constrained and persists throughout the defect’s lifetime.
+Thus, a $^{1}/_{2}$ defect has an offset flow profile throughout with a chirality that is determined at creation (Fig.
+5e).
+The latter drives the bimodal distribution of $^{1}/_{2}$ defect velocities as “up” and “down” defects self-propel with an angular offset (Figs.
+45f, and Extended Fig.
+6.
+This indicates that the chiral symmetry of a $^{1}/_{2}$ is broken at creation and persists until annihilation, where it manifests again in that process.
+During annihilation, the $^{1}/_{2}$ defect tail initially points away from the $^{1}/_{2}$ defect, hence $\varphi^{+}=\pm\pi$.
+In this case a clockwise net torque around the $^{1}/_{2}$ defect results in $\dot{\varphi}^{+}<0$ which causes the defect to rotate into the “up” annihilation configuration $\varphi^{+}>0$ (Fig.
+5b).
+The latter is the chirality associated with defects that originated in the “down” configuration at creation.
+Hence, we expect the chirality of $^{1}/_{2}$ defects created in the “up” configuration to be identical to those that annihilate in the “down” configuration, as indeed observed in experiments (Fig.
+2).
+As in creation, this rotation, combines with defect self-propulsion, results in a curved trajectory (Fig.
+5d).
+In contrast to the creation process, annihilation is independent of $\varphi^{+}$ (assuming $\delta=0$).
+Thus, defects can annihilate even when $\varphi^{+}\neq 0$, as shown in Figs.
+2c-d.
 Time-reversal symmetry breaking
@@
 (2022).
-Here, defects provide physically motivated coarse-graining of the micro-scale motion into a mesoscopically observable feature.
+Here, defects provide physically motivated coarse-graining of the micro-scale motion into a meso-scopically observable feature.
 As a consequence, all the estimates reported here are lower bounds to the actual EP.
@@
 Next, we turn to analyze creation and annihilation events.
-Figures 2c-d show that the distribution of $^{1}/_{2}$ is different close to creation compared to annihilation.
+Figures 2c-d show that the distribution of $\varphi^{+}$ is different close to creation compared to annihilation.
 Similar broken symmetries can also be observed in the average trajectories (Figs.
@@
 1 and 2).
-To quantify this, Table 2 details the results for EP estimates.
-The dimension $D$ indicated the length of the trajectories used (starting at the event).
-Indeed, our results show positive EP both for bacteria and cells.
+Focusing on time-reversal, running the experimental movies backward, creation will appear as annihilation and vice versa.
+To this end, Table 2, column labeled cross, shows the symmetric KL divergence (Jeffreys divergence) between forward-creation/backward-annihilation and forward-annihilation/backward-creation,
+${\color[rgb]{0,0,0}\frac{1}{2}D_{\rm KL}(\overrightarrow{\rm creation}||}$ ${\color[rgb]{0,0,0}\overleftarrow{\rm annihilation})+}$ (1)
+${\color[rgb]{0,0,0}\frac{1}{2}D_{\rm KL}(\overrightarrow{\rm annihilation}||\overleftarrow{\rm creation})}.$
+The positive symmetric divergence indicated that creation and annihilation are not time-symmetric, both for bacteria and cells.
+Comparing the total EP rates, we need to consider the rate of creation and annihilation events.
+For bacteria, the time between events is about 1 sec (50 frames), hence the total EP due to movement is about 50 times larger than due creation-annihilation.
+For cells, the time between creation and annihilation is about 2 hrs (24 frames).
+Hence, EP due to movement is comparable to EP due to creation-annihilation.
+These differences are consistent with the observation that in bacteria, $^{1}/_{2}$ defects appear self-propelled in the sense that their trajectories are ballistic over short time periods Yashunsky et al.
+(2024), implying a relatively large EP.
+With cells, both $\pm^{1}/_{2}$ defects are diffusive (or the timescale of ballistic motion is relatively much shorter) Blanch-Mercader et al.
+(2018), their movement in the bulk is close to reversible, and EP is most apparent in creation and annihilation.
 Table 1: Entropy production (EP) estimation for isolated defects.
@@
 See the methods section for details.
-*Note that the KLD is a non-negative quantity.
-However, the estimator are random variable with an error that is normally distributed with the noted standard deviation.
-Hence, zero EP my result in a negative estimate that is several times the standard error.
-In statistical terms, the hypothesis that $EP=0$ cannot be negated.
 Bacteria Cells
-Event creation annihilation cross creation annihilation cross
-$EP(D=2)$ $075\pm 0.008 0.0$$4\pm 0.008 0.0$$01\pm 0.005^{*}$ $0.63\pm 0.03$ $0.55\pm 0.03$ $0.4\pm 0.3$
-$EP(D=5)$ $0.11\pm 0.01$ $14\pm 0.02 0.0$$06\pm 0.02$ $0.46\pm 0.04$ $47\pm 0.05 0.0$$04\pm 0.03^{*}$
-Focusing on time-reversal, running the experimental movies backward, creation will appear as annihilation and vice versa.
-To this end, Table 2 shows the cross entropy between forward-creation/backward-annihilation and forward-annihilation/backward-creation.
-The dimension $D$ indicated the length of the trajectories used (starting at the event).
-The positive cross-entropy indicated that creation and annihilation are not time-symmetric, both for bacteria and cells.
-Comparing the total EP rates, we need to consider the rate of creation and annihilation events.
-For bacteria, the time between events is about 1 sec (50 frames), hence the total EP due to movement is about 50 times larger than due creation-annihilation.
-For cells, the time between creation and annihilation is about 2 hrs (24 frames).
-Hence, EP due to movement is comparable to EP due to creation-annihilation.
-These differences are consistent with the observation that in bacteria, $^{1}/_{2}$ defects appear self-propelled in the sense that their trajectories are ballistic over short time periods Yashunsky et al.
-(2024), implying a relatively large EP.
-With cells, both $\pm^{1}/_{2}$ defects are diffusive Blanch-Mercader et al.
-(2018), their movement in the bulk is close to reversible, and EP is most apparent in creation and annihilation.
+$0.05\pm 0.01$ $0.5\pm 0.2$
 Conclusions
-Living systems constantly produce and consume energy, keeping them far from thermodynamic equilibrium.
-This continuous energy input drives self-organization, often breaking spatial symmetries in ways that shape the biological function Bowick et al.
-(2022); Pisegna et al.
-(2024); Ardaševa et al.
-(2025).
-A compelling hypothesis suggests that living systems leverage topological mechanisms to regulate their organization, providing a unifying framework for understanding emergent behaviors across biological systems Shankar et al.
-(2022); Andersen et al.
-(2025).
-Here, we explore this idea in two seemingly different systems: Bacillus subtilis bacterial swarms and human bronchial epithelial cells.
-Despite stark differences in biological function and physical properties, both systems exhibit active nematic organization, in which rod-shaped bacteria and elongated polygonal cells align and form dynamic topological defects.
-While bacteria propel in a thin fluid layer at low Reynolds numbers ($\sim 10^{-4}$) Be’er and Ariel (2019); Aranson (2022), HBECs move via substrate adhesion and traction forces.
-In both cases, motion is highly dissipative and sustained only through continuous energy consumption at the cellular level Vaidžiulytė et al.
-(2022).
-Both systems reveal nematic order, where $\pm^{1}/_{2}$ defects undergo continuing creation and annihilation events between defects pairs with opposite charge.
+By studying defect dynamics in two distinct active living systems—swarming Bacillus subtilis bacteria and human bronchial epithelial cell (HBEC) monolayers—we demonstrate that topological defect creation and annihilation are fundamental, symmetry-breaking processes that drive these systems far from equilibrium.
+Despite orders-of-magnitude differences in timescales and propulsion mechanisms, both systems exhibit nematic half-integer defects whose pair dynamics break both mirror symmetry and time-reversal symmetry.
+Moreover, $^{1}/_{2}$ defects spontaneously adopt persistent “up” or “down” chiral configurations and follow spiral trajectories.
+Irreversibility and Entropy Production
 The non-equilibrium and irreversible nature of active living systems is unequivocal.
-Because bacteria and cells are self propelled, the dynamics is out of equilibrium even at the level of the individual.
+Because bacteria and cells are self-propelled, the dynamics is out of equilibrium even at the level of the individual.
 Therefore, the theoretical prediction is that irreversibility should be observed at the single-cell level.
-However, the difficulty in capturing irreversibility from individual positions and orientations is reasonable.
-For any observed single-cell trajectory, one can imagine a cell following the reversed path.
-Thus, EP from single cells is only expected to be captured by adding intra-cellular degrees of freedom corresponding to chemical and biological function, or possibly also the underlying self-propulsion mechanism.
-We can think of defects as physically intuitive, virtual meso-scopic degrees of freedom.
+However, the difficulty in capturing irreversibility from individual trajectories is intuitively clear.
+For any observed single-cell trajectory, one can imagine a cell tracing the reversed path.
+Thus, EP from single cells is only expected to be captured by adding intra-cellular degrees of freedom corresponding to chemical and biological function or the underlying self-propulsion mechanism.
+From this respect, defects are physically intuitive, virtual meso-scopic degrees of freedom.
 Virtual in the sense that they are not the particles of which the system is made of.
-Meso-scopic since each one is defined by a small local collective on several close cells.
-Our findings show that irreversibility is statistically evident locally at the level of the collective.
-Moreover, we find that defect creation and annihilation are not reciprocal processes that general entropy.
-This is also physically intuitive as continuous creation of defects in nematic systems is a hallmark of activity.
-In both experimental systems, entropy production during defect creation and annihilation accounts for a disproportionately large fraction of total entropy production.
+Meso-scopic since each one is defined by a small local collective of several close cells.
+Our findings thus show that irreversibility is statistically evident locally at the level of the collective; see also Tan et al.
+(2019) for measurements of topological entropy and Radhakrishnan et al.
+(2025) for a theoretical study.
+Moreover, we find that defect creation and annihilation are not reciprocal processes and that they generate entropy.
+This is also physically intuitive as the continuous creation of defects in nematic systems is a hallmark of activity.
+In both experimental systems, entropy production during defect creation and annihilation accounts for a disproportionately large fraction of the total entropy production.
 In cells, it is an order of magnitude higher, suggesting that these events are key drivers of non-equilibrium dynamics rather than incidental fluctuations.
-Still, one should remember that our estimates are merely lower bounds for the actual EP, which means that quantitative comparison is problematic.
-We find that mirror symmetry, predicted by linear nematodynamic theory, spontaneously break in the process of creation and annihilation.
-Specifically, when close to each other, $^{1}/_{2}$ spontaneously align perpendicularly (“up” or “down”) to the pair axis (Fig.
-2 e-g).
-Each of these configurations follows spiral-like trajectories rotating either in the clockwise (“up”) or counterclockwise (“down”) direction (Fig.
-3 c-f), further indicating spontaneous breaking of the mirror symmetry.
-We propose that directional self-propulsion disrupts the nematic symmetry that otherwise captures the coarse-grained behavior of bacterial swarms and epithelial cell monolayers.
-This reinforces the role of topological defects in understanding principles of self-organization in living systems.
-Our results challenge the standard active nematic description often applied to active living matter.
-By uncovering these universal physical principles, we take a step toward a deeper understanding of biological self-organization, revealing the fundamental role of symmetry in shaping life at multiple scales.
+Still, one should remember that our estimates are merely lower bounds for the actual EP, which means that quantitative comparison is still lacking.
+Overall, entropy production estimates bring forth some fundamental differences between swarming bacteria and crawling cells regarding the main sources of entropy production (defect motion in bacteria and creation-annihilation in cells).
+This points to the physical mechanisms underlying departure from equilibrium.
+Spontaneous Symmetry Breaking and the Polar-Nematic Model
+Our experiments reveal that topological defect creation and annihilation in living active nematics spontaneously break mirror symmetry—a phenomenon absent in standard active nematic theory.
+Specifically, $^{1}/_{2}$ defects near a $^{1}/_{2}$ defect adopt one of two distinct chiral configurations, exhibiting bimodal orientation distributions and persistent spiral trajectories.
+This chiral symmetry breaking is established during defect creation and persists until annihilation.
+To explain these observations, we introduce an active nemato-polar model that couples a polar vector field $\bm{p}$—representing head–tail asymmetric self-propulsion—to the nematic tensor order parameter $\bm{Q}$.
+A half-integer topological defect in the nematic field imposes a $\pi$-discontinuity (grain boundary) in the polar field, breaking the mirror symmetry of the $^{1}/_{2}$ defect core.
+This generates a net active torque that rotates the defect as it self-propels.
+The torque sign, selected spontaneously during nucleation, determines the chirality and drives the spiral motion of defect pairs.
+The model reproduces key experimental signatures: the bimodal distribution of orientation angle $\varphi^{+}$, the chiral rotation quantified by $\Delta\gamma$, and the flow field offset around $^{1}/_{2}$ defects.
+It also explains the stark contrast with achiral active nematics (e.g., microtubule–kinesin suspensions), where the absence of a polar field yields dynamics consistent with standard nematic theory.
+We suggest that spontaneous symmetry breaking in living active matter stems directly from the intrinsic polar activity of its constituents.
+A complete description therefore requires accounting for the interplay between nematic alignment and polar self-propulsion, revealing how microscopic asymmetry is imprinted onto meso-scopic topological degrees of freedom.
+Overall, our results challenge the standard active nematic description often applied to living active matter.
+By uncovering these universal physical principles, we advance understanding of biological self-organization, revealing the fundamental role of symmetry in shaping life at multiple scales.
 III Methods
@@
 Image analysis was conducted using custom Python code.
-To obtain the velocity field from microscopic images, we employed the Gunnar Farneback method available from OpenCV library producing displacement vectors for every pixel rather than a sparse set of points.
+To obtain the velocity field from microscopic images, we employed the Gunnar Farneback method available from OpenCV library producing displacement vectors for every pixel rather than a sparse set of points Farnebäck (2003).
 The director field was computed using the second-moment matrix, implemented with the structure tensor function from the skimage Python library.
@@
 Defect Detection and Tracking
-Nematic defects were identified using the winding number method, which quantifies the degree of rotation in the nematic field around any closed loop of nearest neighbors.
-When the loop encircles a defect, the winding number reflects the charge of that defect.
-The orientation of the defects was calculated following the method outlined by Vromans & Giomi (2016) Vromans and Giomi (2016).
-Angles are discrete, at integer multiples of $\pi/50$.
-Defect trajectories were automatically tracked using the TrackMate plugin in ImageJ software, which was combined with defect orientation data.
-This generated a list of defects along with their charges, angles, and locations.
+Topological defects were identified by evaluating the winding of the director field.
+We computed the topological charge $k$ at each grid point by summing the angular differences of the director along a closed loop around a $2\times 2$ unit cell.
+Because the nematic phase is apolar, all angular differences were wrapped to the interval $(-\pi/2,\pi/2]$.
+Grid points with a net winding of $\pm\pi$ were labeled as defects with strength $k=\pm^{1}/_{2}$.
+Defect orientations were determined using a direct interpolation method.
+The local director field was interpolated onto a circular path centered on each defect.
+We then extracted the orientation angle by maximizing the agreement between this interpolated profile and the analytical director field for a defect of strength $k$.
+For $^{1}/_{2}$ defects, this yielded a single orientation axis that points toward the comet head.
+For $^{1}/_{2}$ defects, the method resolved the three-fold symmetry of the core and returned the three equivalent orientation angles separated by $120^{\circ}$.
+Defect trajectories were tracked using the trackpy Python library Crocker and Grier (1996).
+The tracking results were merged with the orientation data to produce a list of defects with their charges, orientations, and positions.
 Overall, the data on a single $^{1}/_{2}$ defect trajectory is a sequence of vectors $\textbf{p}_{0},...,\textbf{p}_{K}$, where each vector pi is in $R^{3}$, i.e., each frame consists of 3 scalar values, corresponding to the xyx-y coordinates and an angle, corresponding to the orientation of the $^{1}/_{2}$ defect.
@@
 Quantifying irreversibility in single-defect trajectories
-Here, we provide details for the statistical estimates presented in table 3.
+Here, we provide details for the statistical estimates presented in table  3.
 For each $^{1}/_{2}$ defect, the detection algorithm described above provides a sequences of positions in 2D and orientation angle, $\theta\in[0,2\pi)$.
@@
 For each $^{1}/_{2}$ defect, the detection algorithm described above provides a sequences of positions in 2D and three orientation angles, corresponding to the directions of the three legs $\theta_{1},\theta_{2},\theta_{3}\in[0,2\pi)$.
-We define the average defect direction (taking into account the rank three tensor symmetry) as $p=\sum_{i}(cos\theta_{i},\sin\theta_{i})$.
+We define the average defect direction (taking into account the rank three tensor symmetry) as $p=\sum_{i}(cos3\theta_{i},\sin 3\theta_{i})$.
 The angle corresponding to the defect orientation is then $\varphi=\tan^{-1}(p_{y}/p_{x})/3$.
@@
 The table shows the mean over all defects $\pm$ the standard error.
+The informatic entropy production is the Kullback-Leibler divergence between the distribution of forward observations $\textbf{f}_{i}$ and backward ones $\textbf{b}_{i}$.
+Hence, our experimental data provides us with samples from both spaces.
+In backward trajectories, the sequence of positions is inverted, and the velocity changes sign (because velocities are a forward difference rule).
+However, defect angles do not change; hence the projection also changes sign, hence the distribution of forward projections and backward ones differ.
+Cutting trajectories into 2-frame snippets, we obtain a coarse-grained picture, which is a lower bound of the actual EP.
 […]
@@
 Quantifying irreversibility in defect pairs
-First, the data for each defect trajectory or defect-pair was reduced to include only a few relative angles, as described in the results section.
-Looking only at the first (last) $D$ samples right after (before) a creation (annihilation) event, each event corresponds to an effective vector $\textbf{f}_{i}$.
-For example, looking only at the relative angle between the $^{1}/_{2}$ and $^{1}/_{2}$ defects, each creation event is reduced to a $D$-dimensional vector $\textbf{f}_{i}=(\Theta_{1}\dots\Theta_{D})$, where $\Theta_{1}$ is the angle closest to creation.
-The time-reversed trajectory is simply $\textbf{b}_{i}=(\Theta_{D}\dots\Theta_{1})$.
-The informatic entropy production is the Kullback-Leibler divergence between the distribution of forward observations $\textbf{f}_{i}$ and backward ones $\textbf{b}_{i}$.
-Hence, our experimental data provides us with samples from both spaces.
-In order to evaluate EP, we look at the sequence of projections and find the KLD between the density of forward and backward trajectories.
-In backward trajectories, the sequence of positions is inverted, and the velocity changes sign.
-Defect angles do not change; hence the projection also changes sign.
-Cutting trajectories into 2-frame snippets, we obtain a coarse-grained picture, which is a lower bound of the actual EP.
-Finally, cross entropy between the annihilation and creation is defined as the sum of KLD(forward creation $||$ backward annihilation) + KLD(forward annihilation $||$ backward creation ).
+There are 4 angles associated with the creation/annihilation processes: there related to the defect orientation - $\varphi^{+}$, $\varphi^{-}$ and $\delta$ - and one related to the line connecting the defect centers, $\gamma$.
+We have defined $\gamma$ relative to the angle between the $\pm^{1}/_{2}$ defects at the time of the event.
+Hence, it is not a state function (it cannot be obtained from a single frame).
+Hence, using it may introduce irreversibility, because its measurement is with respect to a fixed event which is either set in the past or future.
+Similarly, $\varphi^{+}$ and $\varphi^{-}$ are relative to $\gamma$.
+However, both $\delta$ and $\Delta\varphi=\varphi^{+}-\varphi^{-}$ are not (here, $\varphi^{-}$ is the angle of the leg which is closes to the comet-tail of the +).
+Therefore, one should use these angles when looking for irreversibility.
+At each event, creation or annihilation, $\delta$ approaches zero.
+Hence, this sense, creation and annihilation are time-symmetric.
+As we demonstrate below, the statistics of how $\delta\to 0$ may be more subtle.
+However, substituting $\delta=0$, $\Delta\varphi=2\varphi^{+}/3+\pi/3$, which is not expected to be reversible.
+More specifically, for each event we look at a short sequence of $D$ snapshots (before annihilation or after creation).
+Looking at $\delta$ and $\Delta\varphi$, a forward trajectory yield a $2D$ dimensional vector $\textbf{f}_{i}=(\delta_{1},\Delta\varphi_{1},\dots,\delta_{D},\Delta\varphi_{D})$.
+The time-reversed trajectory is then $\textbf{b}_{i}=(\delta_{D},\Delta\varphi_{D},\dots,\delta_{1},\Delta\varphi_{1})$.
+Note that it is imperative that the order of $\delta$ and $\Delta\varphi$ does not change.
+Finally, the symmetric KL divergence (Jeffreys divergence) between the annihilation and creation is defined as half the sum of KLD(forward creation $||$ backward annihilation) + KLD(forward annihilation $||$ backward creation ).
 See below for the estimation method.
+The results may depend on the choice of $D$.
+On the one hand, smaller $D$ implies more coarse graining, resulting in a looser lower bound.
+On the other hand, higher $D$ implies a higher dimension for the estimate of the divergence, usually leading to larger statistical errors.
+Since, for any $D$, the estimate provides a lower bound for the actual EP, We may optimize its value.
+We use $D=4$ for bacteria and $D=11$ for cells.
 Estimation of Kullback-Leibler divergence
@@
 where std denotes the standard deviation. Once again, approximating the densities using kkNN yield an estimator for the standard error Beirlant et al. (1997).
-The parameter $k$, setting the number of neighbors, should be taken to be proportional to $\sqrt{N}$ Beirlant et al.
-(1997).
-Estimates reported for single-defects, single cells and the nematic/flow fields used $k=\sqrt{N}$.
-EP of defect pairs use only the orientation of defect, which are discretized.
-As a result, angles are not unique and the distence between neighbors may be zero.
-To this end, we use a larger value of k=2sqrtNk=2sqrt{N}.
-Combining polar and nematic vector fields
-The cells are long, thin objects with a shape that is broadly identical under $\pi$ rotation of the director.
-At high density, they form a thin layer with local orientational order.
-To describe the average orientation local of the we cells we introduce the director field $\hat{n}$ defined as:
-$\underline{\hat{n}}=[\cos(\psi),\sin(\psi)]$ (7)
-where $\psi$ is the average orientation of the cells at a point in space.
-Most observations on the cells are compatible with the typical nematic active stress:
-$\sigma_{ij}\sim(\hat{n}_{i}\hat{n}_{j}-\delta_{ij}/2)$ (8)
-This is the lowest order active term consistent with conservation of momentum and nematic symmetry.
-However, the cells here are able to self-propel, exchanging momentum with the substrate leading to an additional active body force on the material given.
-We introduce the second director field $\hat{p}$, which describes the self-propulsion direction of the cells.
-This can lead to a body force of the form:
-$\underline{f}\sim\underline{\hat{p}}$ (9)
-The cells are elongated in shape and self-propel along their long axis, meaning that $\underline{\hat{p}}$ and $\underline{\hat{n}}$ must be either parallel or anti-parallel, i.e.
-$(\underline{\hat{p}}.\underline{\hat{n}})^{2}=1$.
-We introduce a polar director angle $\theta$ such that:
-$\underline{\hat{p}}=[\cos(\theta),\sin(\theta)]$ (10)
-Which implies that either $\theta=\psi$ or $\theta=\psi+\pi$.
+The parameter $k$, setting the number of neighbors, should be taken to be proportional to $\sqrt{N}$ Beirlant et al. (1997). Estimates reported used $k=\sqrt{N}$.
+Simulating an active nemato-polar system
+We start from the standard description of an incompressible two-dimensional active nematic.
+The nematic orientation of the cells is described by the $\bm{Q}$ tensor which combines the cellular orientation $\bm{\hat{n}}$ and the local degree of nematic order $S\in[0,1]$ into a single tensorial order parameter $Q_{ij}=S(n_{i}n_{j}-\delta_{ij}/2)$.
+The motion of the cells are described by the velocity field $\bm{v}$ which abides by $\nabla.\bm{v}=0$.
+The dynamics of the nematic tensor are given by
+$D_{t}\bm{Q}=\lambda_{Q}S\bm{u}+\frac{1}{\gamma_{Q}}\bm{H}_{Q}+\bm{Q}.\bm{\omega}-\bm{\omega}.\bm{Q},$ (8)
+where $D_{t}=\partial_{t}+\bm{v}.\nabla$ is the material derivative.
+$\bm{u}$ and $\bm{\omega}$ represent the strain rate and vorticity tensors, respectively.
+The molecular tensor is given by the functional derivative of the elastic free energy with respect to the nematic tensor, $\bm{H}=-\delta F/\delta\bm{Q}$ and $\gamma_{Q}$ is the rotational viscosity associated with the nematic field.
+The variable $\lambda_{Q}$ describes the tendency of the nematic field to align with the flow.
+To describe the polar orientation of the cells we introduce the vectorial order parameter $\bm{p}$ which describes the average local orientation of the self-propulsion direction of the cells.
+This follows similar dynamics to the nematic tensor, given by
+$D_{t}\bm{p}=\lambda_{p}\bm{u}.\bm{p}+\frac{1}{\gamma_{p}}\bm{h}_{p}-\bm{\omega}.\bm{p}.$ (9)
+Here, the molecular vector is given by the functional derivative of the elastic free energy with respect to the polar field, $\bm{h}=-\delta F/\delta\bm{p}$ and $\gamma_{p}$ is the rotational viscosity associated with the polar field.
+As before, the variable $\lambda_{p}$ describes the tendency of the polar field to align with the flow.
+The elastic free energy governs the relaxation dynamics of the $\bm{Q}$ and $\bm{p}$ fields. We split it into 3 parts,
+$F=F_{Q}+F_{C}+F_{p}.$ (10)
+The first part contains the nematic energy and describes the steric interactions between cells. It is given by
+$F_{Q}=\frac{1}{2}\int\Big[K_{Q}|\nabla\bm{Q}|^{2}+\alpha_{Q}\textrm{tr}\bm{Q}^{2}(\textrm{tr}\bm{Q}^{2}-1)\Big]\textrm{d}A.$ (11)
+In the expression above, $K_{Q}$ denotes the single elastic constant and $\alpha_{Q}$ controls the strength of the Landau terms which apply a soft constraint pushing the system into the nematic phase $S=1$.
+The second part of the energy is the coupling term between the $\bm{Q}$ and $\bm{p}$ fields given by
+$F_{C}=-\frac{\alpha_{C}}{2}\bm{p}.\bm{Q}.\bm{p}.$ (12)
+This energy is minimized when the $\bm{Q}$ and $\bm{p}$ fields align and is used to ensure they remain compatible descriptions of the cellular orientations.
+The final part of the energy describes the tendency to locally align polarity and is given by
+$F_{p}=\frac{1}{2}\int\Big[K_{p}|\nabla\bm{p}|^{2}+\frac{\alpha_{p}}{4}\bm{p}^{4}-\frac{\beta_{p}}{2}\bm{p}^{2}\textrm{d}A.$ (13)
+Here, $K_{p}$ is the elastic constant associated with gradients in the polar field and $\alpha_{p}$ and $\beta_{p}$ control the transition to polar order.
+This energy is employed predominantly to maintain the polar order parameter and keep the field continuous, thus it has a smaller magnitude than the other energies.
+We typically set $\beta_{p}=0$.
+This implies that the polar field becomes disordered when there is no nematic order, for example at the core of a defect.
+The dynamics of the velocity evolve according to
+$\rho D_{t}\bm{v}=\eta\nabla^{2}\bm{v}-P\bm{I}+\nabla.\bm{\sigma}^{p}+\zeta_{Q}\nabla.\bm{Q}+\zeta_{p}\bm{p}-\mu\bm{v}.$ (14)
+Here, $\rho$ is a constant density, $\eta$ is the shear viscosity, and $\mu$ is the friction coefficient.
+$\zeta_{Q}$ is the magnitude of the nematic active stress and $\zeta_{p}$ is the magnitude of the polar active force.
+$P$ if the pressure which is fixed by the incompressibility condition and $\bm{I}$ is the identity matrix.
+Finally, the passive stress is given by
+$\bm{\sigma}^{p}=-\lambda_{Q}S\bm{H}+\bm{Q}.\bm{H}-\bm{H}.\bm{Q}-\lambda_{p}\bm{h}\odot\bm{p}-\bm{h}\wedge\bm{p}.$ (15)
+The symmetric and antisymmetric tensor products are defined as
+$\bm{h}\odot\bm{p}$ $=\frac{1}{2}(h_{i}p_{j}+h_{j}p_{i})$ (16)
+$\bm{h}\wedge\bm{p}$ $=\frac{1}{2}(h_{i}p_{j}-h_{j}p_{i}).$ (17)
+We simulate the dynamical equations on a $512\times 512$ descretized grid using finite differences to approximate the gradients.
+We use a vorticity stream function approach to solve for incompressibility.
+We non-dimensionalize our equations using the length scale $\epsilon=\sqrt{K_{Q}/\alpha_{Q}}=1$, and time scale $\tau=\gamma_{Q}/\alpha_{Q}=1$, setting our simulation size to $L=128$.
+We measure energy in units of the elastic constant, $K_{Q}=1$.
+The remaining parameters are given by $\rho=1$, $\eta=1$ and $\mu=16$.
+We enforce low Reynolds number by neglecting the advective term in Eq.
+14.
+We simulate the system in the absence of flow alignment, $\lambda_{Q}=\lambda_{p}=0$.
+To ensure the polar field follows the nematic field closely, we set $K_{p}=K_{Q}/100$, $\alpha_{p}=\alpha_{Q}/25$ and $\alpha_{C}=4\alpha_{Q}/5$.
+Finally we fix the nematic active stress at $\zeta_{Q}=-0.2$ and the polar active force is taken from the range $\zeta_{P}\in[0.0,0.0026125]$.
+We use this model to simulate specific creation and annihilation events, as this allows us to keep the grain boundary correlated with the orientation of the $^{1}/_{2}$ defect.
+The system is initialized with zero flow, $\underline{v}=0$.
+To simulate defect creation, we initialize the director fields with a localized bend distortion, see Extended Data Fig.
+7a.
+If $\theta$ is the orientation of the polar field the bend distortion is given by
+$\theta(\underline{r})=\frac{\pi}{2}+\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)\sin\left(\frac{2\pi r_{y}}{L}\right)$ (18)
+Here $\underline{r}=(r_{x},r_{y})$ is the position relative to the center of the simulation area of size $L$ and $\sigma=11L/128$ gives the scale of the distortion.
+For defect annihilation, we initialize the system with a pair of defects situated either side of the center of the simulation area, see Extended Data Fig.
+7b.
+The field around a defect is given by
+$\theta(\underline{r})=\frac{\phi_{1}-\phi_{2}}{2}.$ (19)
+Here $\phi_{1}$ and $\phi_{2}$ are the polar angles around the locations of the $^{1}/_{2}$ and $^{1}/_{2}$ defects, respectively.
+The defects are located at $\underline{r}_{\mp}=(\pm\Delta,0)$ and $\Delta$ is half the separation between the defects; we set $\Delta=3L/32$.
+Finally, we reverse all polar vectors above the $x$ axis, which moves the grain boundary to the tail of the $^{1}/_{2}$ defect.
 Field around a topological defect
-Topological defects within the nematic field are characterized by a half integer winding number $k=\pm^{1}/_{2}$. The director field at a polar angle $\phi$ around a topological defect is given by:
+As above we describe the nematic field by the tensor $\bm{Q}=S(\bm{n}\odot\bm{n}-\bm{I}/2)$ where $\bm{n}$ is a vector aligned with the nematic director.
+We now introduce the nematic orientation angle, $\psi$ which describes the average orientation of the nematic director at a point in space, hence
+$\hat{\bm{n}}=[\cos(\psi),\sin(\psi)].$ (20)
+Topological defects within the nematic field are characterized by a half integer winding number $k=\pm^{1}/_{2}$.
+The orientation of the nematic director field at a polar angle $\phi$ around a topological defect is given by:
 $\psi=k\phi+\psi_{0}$ (21)
@@
 3.
-A half integer winding number is incompatible with the polar field $\hat{p}$, and necessitates the introduction of a discontinuous jump in the polar director angle, $\theta$.
-This is the grain boundary over which $\underline{\hat{p}}$ goes through a $\pi$ rotation which begins and ends at a topological defect.
+We describe the polar field with $\bm{p}$ and introduce the polar orientation angle, $\theta$, which is compatible with
+$\hat{\bm{p}}=p[\cos(\theta),\sin(\theta)].$ (22)
+For the following calculations we will neglect variations in the order parameter, i.e. $S=p=1$ and assume perfect correlation between the polar and nematic fields, i.e. $\bm{p}.\bm{n}=\pm 1$.
+A half integer winding number, $k=\pm^{1}/_{2}$, is incompatible with the polar field $\bm{p}$, and necessitates the introduction of a discontinuous jump in the polar director angle, $\theta$.
+This is the grain boundary over which $\bm{{p}}$ goes through a $\pi$ rotation which begins and ends at a topological defect.
 We introduce the grain boundary as follows
@@
 3.
+Flow field around a topological defect
+At low Reynolds number, the flow caused by a force field $\bm{f}$ obeys the following equation
+$\eta\nabla^{2}\bm{v}-\nabla P+\bm{f}=0.$ (25)
+Where $\eta$ is the viscosity and $P$ is the pressure.
+For an incompressible fluid, solutions to this equation can be found by convoluting the force field $\bm{f}$ with the two dimensional Oseen tensor $\bm{G}$, hence
+$v_{i}(\bm{r})=\int G_{ij}(\bm{r}-\bm{r}^{\prime})f_{j}(\bm{r}^{\prime})dA^{\prime},$ (26)
+where the Oseen tensor is given by
+$G_{ij}(\bm{r})=\frac{1}{4\pi\eta}\left[(\log{\frac{\mathcal{L}}{{r}}}-1)\delta_{ij}+\frac{r_{i}r_{j}}{r^{2}}\right]$ (27)
+To calculate the flow around a defect, we assume that the active forces dominate, neglecting the passive stress tensor.
+In our model, the flow around a defect is dominated by the two active forces.
+The active force arising from the extensile active stress tensor is given by
+$\bm{f}_{Q}=\zeta_{Q}\nabla.\bm{Q},$ (28)
+and can be calculated as
+$\bm{f}_{Q}=\frac{\zeta_{Q}}{2r}\bm{\hat{x}}$ (29)
+where $r$ is the distance from the defect.
+Following the calculation in Giomi et al. (2014) one can arrive at the following flow field arising from the active nematic forces
+$\bm{v}^{Q}(r,\phi)=\frac{\zeta_{Q}}{12\eta}\left[3(R-r)+r\cos(2\phi),r\sin(2\phi)\right].$ (30)
+Due to the linearity of Eq. 25, we can simply add this to the additional velocity caused by the polar body force.
+The polar body force arising from self-propulsion is given by
+$\bm{f}_{p}=\zeta_{p}\bm{p}.$ (31)
+This can be calculated around a defect as
+$\bm{f}_{p}=\zeta_{p}(-1)^{m}S(\phi,\phi_{g})\cos\left(\frac{\phi}{2}\right)\hat{\mathbf{x}}+\zeta_{p}(-1)^{m}S(\phi,\phi_{g})\sin\left(\frac{\phi}{2}\right)\hat{\mathbf{y}}$. (32)
+Here we have introduced the sign function defined by
+[…]
+Following a similar process to that shown in Ref Giomi et al.
+(2014), one can arrive at a flow field that results from the active polar body force, which we denote $\bm{v}^{p}$.
+The equation representing $\bm{v}^{p}$ is very long and a full derivation is given in the appendix, along with the equivalent calculation for a $^{1}/_{2}$ defect.
+The final flow field around an active $^{1}/_{2}$ defect is given by $\bm{v}=\bm{v}^{Q}+\bm{v}^{p}$ and is shown in Fig.
+5e.
+To assess the active self-propulsion of a $^{1}/_{2}$ defect, we evaluate the generated velocity at the defect core, i.e. at $r=0$. This gives
+$v(r=0)=\frac{R\zeta_{Q}}{4\pi\eta}\left[\pi+\frac{\zeta_{p}}{\zeta_{Q}}R\sin\left(\frac{\phi_{G}}{2}\right),-\frac{\zeta_{p}}{\zeta_{Q}}R\cos\left(\frac{\phi_{G}}{2}\right)\right].$ (34)
+The argument of this velocity vector is displayed in Fig. 5f (evaluated for $R=1$ and $\eta=1$).
+The torque on the defect is estimated from the curl of the velocity field at $r=0$, which is calculated numerically and displayed in Fig. 5g.
 Total force on a topological defect core due to polar forcing
+Calculating the flow field around the defect is often a difficult process, however calculating the net force at the defect core is comparatively easy and gives very similar results.
+We include it here for completeness.
 We now calculate the total additional force on the defect core due to polar forcing by integrating the polar force azimuthally around the core of the defect.
@@
 The total body force around the core of the defect with charge $k$ is given by.
-$\underline{F}$ $=\oint\underline{f}(\phi)\textrm{d}\phi$ (14)
-$\sim\oint\underline{\hat{p}}(\phi)\textrm{d}\phi$ (15)
+$\underline{F}_{p}$ $=\oint\underline{f}_{p}(\phi)\textrm{d}\phi$ (35)
+$\sim\zeta_{p}\oint\underline{\hat{p}}(\phi)\textrm{d}\phi$ (36)
 It is convenient to evaluate this integral piecewise to find the solution:
-$\underline{F}$ $\sim\int_{-\pi}^{\phi_{g}}\underline{\hat{p}}(\phi)\textrm{d}\phi+\int_{\phi_{g}}^{\pi}\underline{\hat{p}}(\phi)\textrm{d}\phi$
-$=\int_{-\pi}^{\phi_{g}}[\cos(k\phi+m\pi),\sin(k\phi+m\pi)]\textrm{d}\phi+$
-$+\int_{\phi_{g}}^{\pi}[\cos(k\phi+(m+1)\pi),\sin(k\phi+(m+1)\pi)]\textrm{d}\phi$
-$=\left[\frac{[\sin(k\phi+m\pi),-\cos(k\phi+m\pi)]}{k}\right]_{-\pi}^{\phi_{g}}+$
-$+\left[\frac{[\sin(k\phi+(m+1)\pi),-\cos(k\phi+(m+1)\pi)]}{k}\right]_{\phi_{g}}^{\pi}$
-$=\frac{2}{k}\left[\sin\left(k\phi_{g}+m\pi\right),-\cos\left(k\phi_{g}+m\pi\right)\right]$
-$=\frac{2}{k}\left[\cos\left(k\phi_{g}+(m-0.5)\pi\right),\sin\left(k\phi_{g}+(m-0.5)\pi\right)\right]$
+$\underline{F}$ $\sim\zeta_{p}\int_{-\pi}^{\phi_{g}}\underline{\hat{p}}(\phi)\textrm{d}\phi+\zeta_{p}\int_{\phi_{g}}^{\pi}\underline{\hat{p}}(\phi)\textrm{d}\phi$
+$=\zeta_{p}\int_{-\pi}^{\phi_{g}}[\cos(k\phi+m\pi),\sin(k\phi+m\pi)]\textrm{d}\phi+$
+$\zeta_{p}\int_{\phi_{g}}^{\pi}[\cos(k\phi+(m+1)\pi),\sin(k\phi+(m+1)\pi)]\textrm{d}\phi$
+$=\zeta_{p}\left[\frac{[\sin(k\phi+m\pi),-\cos(k\phi+m\pi)]}{k}\right]_{-\pi}^{\phi_{g}}+$
+$\zeta_{p}\left[\frac{[\sin(k\phi+(m+1)\pi),-\cos(k\phi+(m+1)\pi)]}{k}\right]_{\phi_{g}}^{\pi}$
+$=\frac{2\zeta_{p}}{k}\left[\sin\left(k\phi_{g}+m\pi\right),-\cos\left(k\phi_{g}+m\pi\right)\right]$
 This force is shown as the large red arrow in Extended Data Fig. 3. Some important things to note here.
 First, there is a net translational force on both $\pm^{1}/_{2}$ defects, meaning that we expect both defects to self-propel.
-This is in addition to the self-propulsion force already present for $^{1}/_{2}$ defects due to the nematic active stress.
 If $\phi_{g}=0$, as is the case immediately after creation, the additional force is perfectly perpendicular to the symmetry axis of the defect.
-This would imply a self-propulsion perpendicular to the tail for $^{1}/_{2}$ defects (or a leg for $^{1}/_{2}$ defects).
+When combined with the force from the extensile stress, this would imply a self-propulsion at an angle to the tail for $^{1}/_{2}$ defects.
 The additional force on the defects breaks chiral symmetry provided it does not align with a mirror symmetry axis of the defect.
@@
 As mentioned in the main text, $m$ arises from a spontaneously broken symmetry.
+Finally, the force on the core of a $k=+^{1}/_{2}$ defect is perfectly aligned with the flow predicted by the Stokes flow calculation.
+If we add this to the net force arising from the extensile active stress, which is given by $\underline{F}_{Q}\sim[\zeta_{Q},0]$ we arrive at
+$\underline{F}=\left[\zeta_{Q}+\zeta_{p}\sin\left(\phi_{g}/2\right),-\zeta_{p}\cos(\phi_{g}/2)\right],$ (38)
+which shows the same behavior and symmetries as Eq.
+34, see Extended Data Fig.
+3c.
+Thus the self-propelled motion of the $^{1}/_{2}$ defects can be simply understood by considering the linear combination of the active forces at the core.
 Total torque on a topological defect core due to polar forcing
-We can assess the active torsional force around a defect caused by the polar active force by azimuthally integrating the azimuthal component of the polar force. This is given by:
+We can assess the active torsional force around a defect caused by the polar active force by azimuthally integrating the azimuthal component of the polar force.
+This gives an equivalent picture to the predicted vorticity at the core of the defect.
+The curl of the polar force is given by:
 $\tau=\oint\underline{f}.\underline{\textrm{d}l}$ (39)
 This integration is performed on a circle around the defect and $\underline{dl}$ is tangent to that circle. Thus we can write:
-$\tau\sim\int_{-\pi}^{\pi}\underline{\hat{p}}.[-\sin(\phi),\cos(\phi)]\textrm{d}\phi$ (18)
+$\tau\sim\zeta_{p}\int_{-\pi}^{\pi}\underline{\hat{p}}.[-\sin(\phi),\cos(\phi)]\textrm{d}\phi$ (40)
 We can simplify this with angle addition formula to obtain:
-$\tau$ $\sim\int_{-\pi}^{\pi}\sin(\theta-\phi)\textrm{d}\phi$ (19)
-$=\int_{-\pi}^{\pi}\sin((k-1)\phi+\pi*(H(\phi-\phi_{g})+m))\textrm{d}\phi$ (20)
-$=\left[\frac{-\cos((k-1)\phi+m\pi)}{k-1}\right]_{-\pi}^{\phi_{g}}+$ (21)
-$+\left[\frac{-\cos((k-1)\phi+(m+1)\pi)}{k-1}\right]_{\phi_{g}}^{\pi}$ (22)
-$=\frac{1}{k-1}\left[2\cos((k-1)\phi_{g}+(m+1)\pi)+\right$. (23)
+$\tau$ $\sim\zeta_{p}\oint_{-\pi}^{\pi}\sin(\theta-\phi)\textrm{d}\phi$ (41)
+$=\zeta_{p}\oint_{-\pi}^{\pi}\sin((k-1)\phi+\pi*(H(\phi-\phi_{g})+m))\textrm{d}\phi$ (42)
+$=\zeta_{p}\left[\frac{-\cos((k-1)\phi+m\pi)}{k-1}\right]_{-\pi}^{\phi_{g}}+$ (43)
+$\zeta_{p}\left[\frac{-\cos((k-1)\phi+(m+1)\pi)}{k-1}\right]_{\phi_{g}}^{\pi}$ (44)
+$=\frac{\zeta_{p}}{k-1}\left[2\cos((k-1)\phi_{g}+(m+1)\pi)+\right$. (45)
 $\quad\left.\cos(m\pi)-\cos((2k+m-1)\pi)\right]$ (46)
-$=\frac{2}{1-k}\cos((k-1)\phi_{g}+m\pi)$ (25)
+$=\frac{\zeta_{p}}{1-k}\cos((k-1)\phi_{g}+m\pi)$ (47)
 This torque is shown as the large green arrow in Extended Data Fig. 3.
@@
 The torque disappears when $(k-1)\phi_{g}+m\pi=(n+0.5)\pi$, $n\in\mathbb{Z}$.
-This is equivalent to when the net additional net force aligns with an axis of miror symmetry.
+This is equivalent to when the net additional net force aligns with an axis of mirror symmetry.
 Third, the magnitude of the torque additionally depends on $k$, with $^{1}/_{2}$ defects experiencing three times more torque on their core than $^{1}/2$ defects.
-IV Supplementary material
-Table 3: Entropy production (EP) estimation - control.
-Top: EP for trajectories of individual bacteria and cells.
-Bottom: EP for the nematic and flow fields at random positions.
-Values show the estimate $\pm$ the standard error.
-Since this is a lower bound, a value that is statistically larger than zero implies non reversible dynamics.
-See the methods section for details.
-Bacteria Cells
-EP individual trajectories $004\pm 0.004 0.0$$003\pm 0.0002$
-EP nematic+flow $008\pm 0.001 0.000$$008\pm 0.001$
+Finally, since the curl of the force arising from the active stress is zero around the core of a defect, any net vorticity around the defect must be arise from the curl of the body force.
+We plot $\tau$ around a $^{1}/_{2}$ defect as a function of $\phi_{g}$ and $\zeta_{p}$ in Extended Data Fig.
+3d, and we see that it is a good predictor for the vorticity around a defect (Fig.
+5g.

### v3

@@
 The source of mirror symmetry breaking
-To describe cell self-propulsion, we introduce a polar field, $\underline{p}$, indicating the direction of cell movement relative to the substrate, alongside the nematic field that describes the cells’ long axes.
-This can be visualized as two superimposed parallel fields: one with lines for the long axes and another with arrows for the head direction (see Fig.
-5a).
-For simplicity, we assume that both $\underline{p}$ and $\underline{p}$, the latter pointing in the opposite direction, are compatible with the same nematic field.
-During creation in extensile nematics, defects arise from a bend instability that introduces a net curl to the polar field and breaks chiral symmetry (Fig.
-5a), with a particularly strong net curl around the $^{1}/_{2}$ defect (green arrow; Fig.
-5a).
-Notably, the mirror image of this process produces defects with opposite chirality, indicating that defect chirality selection is driven by spontaneous symmetry breaking.
-In this context, distinct “up” and “down” defect configurations should emerge during creation.
-Around an isolated defect, the polar field exhibits a net polarization and curl that depend on the grain boundary position, see SI for a derivation.
-This polar field indicates the force direction exerted by neighboring particles, generating both force and torque on the defect core (Fig.
-5a).
-A similar symmetry breaking mechanism has been noted in simulations of self-propelling thin rods and flexible polymers Palmer et al.
-(2022); Prathyusha et al.
-(2018); Joshi et al.
-(2019).
-We suggest that defect dynamics and flows around the defects in our experiments result from a combination of torque and extensile active stress.
-For a $^{1}/_{2}$ defect, this leads to an offset force along the defect’s symmetry axis and an additional net torque.
-$^{1}/_{2}$ defect experiences smaller net torques and forces, allowing for self-propulsion and rotation.
-The calculated Stokes flow field around isolated defects (Figs.
-5c-d) indicates that immediately after the defects are created, flow patterns exhibiting opposite chiral symmetry are equally likely.
-This symmetry breaking can explain the distinct behavior of “up” and “down” configurations as follows.
-Defects are generated through a bend instability induced by extensile active stress, hence are generated with $\varphi^{+}\approx 0$.
-After creation, $^{1}/_{2}$ move in a curved trajectory, spiraling away from its $^{1}/_{2}$ pair due to the non-zero curl resulting from the polar field.
-Figure 5a illustrates a situation where a $^{1}/_{2}$ defect moves, rotating clockwise.
-Consequently, this leads to the defect pair to also rotate clockwise, which we refer to as “up” configuration (see Fig.
-4).
-The polar grain boundary connected to the defect core is topologically constrained and persists throughout the defect’s lifetime.
-As a result, the same broken symmetry influences the defect annihilation process, as illustrated in Fig.
-5b.
-When a $^{1}/_{2}$ defect approaches annihilation, two key forces dictate its trajectory.
-The first is the elastic attraction to the $^{1}/_{2}$ defect with which it is annihilating, and the second is the active self-propulsion force.
-The chiral torque acting on the defect causes it to rotate, so it does not self-propel directly toward the $^{1}/_{2}$ defect; however, the elastic forces prevent complete escape.
-Consequently, the resulting trajectory forms an inward spiral toward annihilation, as depicted in Figs.
-3e-f.
-In contrast to the creation process, annihilation is independent of $\varphi^{+}$ (assuming $\delta=0$).
-Thus, defects can annihilate even when $\varphi^{+}\neq 0$, as shown in Figs.
-2c-d.
-Additionally, the reverse of this process is equally likely, resulting in the observed “up’ and “down’ configurations during annihilation.
 In this section we provide a possible physical explanation to the symmetry breaking observed above.